### v2

# Mining Glitch Tokens in Large Language Models via Gradient-based Discrete Optimization

Abstract
Glitch tokens in Large Language Models (LLMs) can trigger unpredictable behaviors, compromising model reliability and safety. Existing detection methods often rely on manual observation to infer the prior distribution of glitch tokens, which is inefficient and lacks adaptability across diverse model architectures. To address these limitations, we introduce GlitchMiner, a gradient-based discrete optimization framework designed for efficient glitch token detection in LLMs. GlitchMiner leverages an entropy-based loss function to quantify the uncertainty in model predictions and integrates first-order Taylor approximation with a local search strategy to effectively explore the token space. Our evaluation across various mainstream LLM architectures demonstrates that GlitchMiner surpasses existing methods in both detection precision and adaptability. In comparison to the previous state-of-the-art, GlitchMiner achieves an average improvement of 19.07% in precision@1000 for glitch token detection. By enabling efficient detection of glitch tokens, GlitchMiner provides a valuable tool for assessing and mitigating potential vulnerabilities in LLMs, contributing to their overall security. Our code is available at https://github.com/wooozihui/GlitchMiner.
1 Introduction
Large Language Models (LLMs), such as GPT-4 gpt4 , have revolutionized natural language processing, enabling breakthroughs in critical fields, including code generation LLM_code1 ; LLM_code2 ; LLM_code3 , medical research medical1 ; medical2 , and education education1 ; education2 . Despite their strengths, these models harbor a subtle yet critical vulnerability: glitch tokens geiping2024coercing . Glitch tokens are anomalous tokens that disrupt the model’s ability to understand or generate accurate responses, leading to unpredictable behaviors such as inappropriate content, repetitive errors, hallucinated outputs, or offensive language geiping2024coercing ; solidgoldmagikarp ; GlitchHunter ; Magikarp . This issue stems from the tokenization process, potentially arising due to insufficient training or poorly embedded tokens within the model’s vocabulary.
For instance, as shown in Figure 1, the Llama2-7B-chat-hf llama2 model fails to recognize the token “oreferrer,” treating it as an uncorrelated term. This oversight reveals deeper vulnerabilities in token handling, leading to unexpected outputs that can compromise the reliability of LLMs, especially in high-stakes applications. Mitigating these glitches is crucial to ensuring the trustworthiness and robustness of LLM-generated content.
Existing methods for detecting glitch tokens primarily rely on manual observation to assume their prior distribution solidgoldmagikarp ; GlitchHunter ; Magikarp . However, this approach is not only time-consuming and labor-intensive but also struggles to adapt to different model architectures and vocabularies. Moreover, it often overlooks glitch tokens that fall outside predefined distribution patterns, limiting the detection’s coverage and accuracy.
To overcome these challenges, we propose GlitchMiner, a novel gradient-based discrete optimization framework for efficiently detecting glitch tokens in LLMs. Unlike previous methods that rely on manual observation of glitch tokens, GlitchMiner focuses on identifying tokens that deviate significantly from the normal token distribution—where most tokens exhibit high certainty and predictable behavior. As illustrated in Figure 2, GlitchMiner leverages the uncertainty in token predictions to uncover potential glitches. It employs an entropy-based loss function to measure uncertainty and uses a local discrete optimization strategy, combining first-order Taylor approximation Taylor with local search, to efficiently explore the token space. This adaptive approach enables GlitchMiner to surpass the limitations of earlier methods, providing more comprehensive detection without relying on predefined distribution assumptions.
Our main contributions in this paper include:
  1. 1.
    Entropy-based Loss Function: We propose using entropy as the optimization objective, effectively capturing the uncertainty in model predictions without relying on manually defined glitch token prior distributions.
  2. 2.
    Local Search Strategy: We design a novel discrete optimization method that combines first-order Taylor approximation with local search, significantly improving the efficiency and accuracy of glitch token detection.
  3. 3.
    Wide Applicability: Through experiments on multiple mainstream LLM architectures, we demonstrate the generalizability and effectiveness of GlitchMiner. Experimental results show that GlitchMiner improves the precision@1000 of glitch token detection by an average of 19.07% compared to previous state-of-the-art method.
  4. 4.
    Efficient Detection: Our method can quickly locate potential glitch tokens in large-scale vocabularies, greatly improving detection efficiency and providing a powerful tool for assessing and improving the security of LLMs.
The structure of this paper is as follows: Section 2 reviews related work, including glitch token detection and gradient-based discrete optimization methods; Section 3 details the design and implementation of GlitchMiner; Section 4 presents experimental results and analysis; finally, Section 5 concludes the paper and discusses future research directions.
Figure 1: Example of a glitch token "oreferrer", causing misinterpretation in the Llama2 model.
Figure 2: Overview of the GlitchMiner Framework. This figure illustrates the two-stage process: (1) Initialization Stage – filters out unnecessary tokens to reduce the search space; (2) Mining Stage – iteratively selects batches of tokens, evaluates their entropy, and determines whether they are glitch tokens.
2 Related Work
2.1 Glitch Token Detection
Detecting glitch tokens in LLMs has garnered increasing attention. Several methods have been proposed to address this issue, including Magikarp Magikarp , GlitchHunter GlitchHunter , and GlitchProber GlitchProber , each utilizing different strategies to identify glitch tokens.
Magikarp Magikarp detects glitch tokens by analyzing token embeddings. It uses characteristic patterns, such as the $\ell_{2}$ norm, in the embedding space to identify under-trained tokens and employs simple repetition tasks to test whether the model can accurately reproduce these tokens. This method focuses on quickly detecting problematic tokens through direct queries, making it fast and lightweight.
GlitchHunter GlitchHunter adopts a clustering-based approach. It assumes that glitch tokens tend to cluster in the embedding space and constructs a Token Embedding Graph (TEG), where nodes represent tokens connected based on their embedding similarity. By applying the Leiden algorithm traag2019louvain , GlitchHunter identifies potential glitch clusters. Hypothesis testing is then used to refine these clusters, iterating multiple times to ensure the accuracy of detection.
GlitchProber GlitchProber takes a different approach by analyzing the internal activations within transformer layers, such as attention heads and hidden states. It reduces the dimensionality of these activations using PCA PCA and applies SVM classifiers SVM to identify glitch tokens. Additionally, it integrates mitigation mechanisms by modifying neuron activations to minimize the impact of glitch tokens during inference.
Both Magikarp and GlitchHunter rely on manually observed prior distributions to sample token space for glitch token detection, performing well when their assumptions align with the actual glitch token distribution. However, their adaptability can be limited when real-world distributions deviate from these priors. GlitchProber aims to improve detection through automation, but it requires a large number of queries to collect sufficient training data. Additionally, the need for extra training introduces computational overhead, impacting efficiency in practice.
Our method overcomes these limitations by using entropy-based loss functions to quantify uncertainty in predictions, without relying on priors or additional training.
2.2 Gradient-based Discrete Optimization
Gradient-based discrete optimization methods hotflip ; autoprompt ; GCG leverage gradient information to predict how individual tokens impact the loss function. These approaches typically treat the one-hot encoding of tokens or token embeddings as continuous vectors to compute gradients, guiding token replacements for optimization.
HotFlip hotflip uses the one-hot encoding of tokens to compute gradients and selects the token with the largest negative gradient to replace the current token, aiming to minimize the loss. However, it only evaluates one candidate token per iteration, which can lead to suboptimal predictions and reduced accuracy.
AutoPrompt autoprompt improves upon HotFlip by evaluating multiple candidate tokens in each iteration. Instead of relying on gradients from one-hot encodings, it utilizes token embedding gradients for loss estimation, enhancing prediction accuracy by considering a broader range of potential token replacements.
GCG GCG extends HotFlip by incorporating multi-candidate token selection, similar to AutoPrompt, but it still uses the one-hot encoding of tokens to compute gradients for loss estimation. Notably, GCG has been applied to automated jailbreaks shen2023anything in LLMs, efficiently searching for adversarial suffixes.
Both AutoPrompt and GCG rely on large batch sampling to reduce inaccuracies in gradient predictions. These inaccuracies occur because Taylor approximation become less reliable when the input tokens are far from the original reference points.
To address this challenge, we introduce a local search strategy in our approach. By focusing on a smaller, localized token space, our method improves precision without the need for large batch sampling. This targeted search not only enhances precision but also avoids unnecessary queries over non-glitch tokens, significantly boosting query efficiency for glitch token detection.
3 Method
Algorithm 1 GlitchMiner Framework
1:Input: Token set $\mathcal{T}$, Iteration number $I$, Batch size $B$
2:Output: Glitch token set $\mathcal{G}$
3:# Stage 1: Initialization
4:$\mathcal{T}^{*}\leftarrow\text{Filter}(\mathcal{T})$ Filter out the unnecessary token set $\mathcal{T}^{*}$
5:$\mathcal{G}\leftarrow\emptyset$
6:# Stage 2: Mining
7:for $i\in I$ do
8:     Select a batch of tokens $\mathcal{B}$ from $\mathcal{T}\setminus(\mathcal{T}^{*}\cup\mathcal{G})$, where $|\mathcal{B}|=B$
9:     for each token $t\in\mathcal{B}$ do
10:         if Verify($t$then
11:              Add $t$ to $\mathcal{G}$
12:         else
13:              Add $t$ to $\mathcal{T}^{*}$
14:         end if
15:     end for
16:end for
GlitchMiner Framework. As shown in Algorithm 1, GlitchMiner operates in two main stages: initialization and mining.
In the initialization stage, we filter out tokens that don’t require detection to reduce the search space and improve efficiency, after which the mining stage iteratively selects batches of unverified tokens for verification. Each token is checked: if it is identified as a glitch token, it is added to the discovered set; otherwise, it is marked unnecessary for further detection. This process continues until a stopping condition is met, such as reaching a specified number of iterations or detected glitch tokens.
The following sections detail how GlitchMiner implements its key components:
  1. 1.
    Token Filter: Identifying tokens that don’t require detection.
  2. 2.
    Glitch Token Verification: Methods to verify if a token is a glitch token.
  3. 3.
    Token Selection Strategy: Selecting the most promising glitch token candidates in each iteration.
By exploring these components, we demonstrate how GlitchMiner efficiently implements glitch token mining in practice.
3.1 Token Filter
In the initialization stage of GlitchMiner, we first filter all tokens to remove those that do not need to be detected. Similar to Magikarp , we designed a token filter module that classifies tokens based on their characteristics and filters out those that are unnecessary for detection.
The core idea of the token filter module is to classify each token by decoding and then re-encoding it, ensuring it meets specific classification criteria. Specifically, we prepend a special prefix "«" to each token to maintain consistency during the encoding and decoding process. Then, we filter out tokens that fall into the following categories:
  • SPECIAL: Special tokens, such as [BOS], </s>, etc., which are enclosed in brackets or angle brackets and contain alphabetic characters.
  • UNDECODEABLE: Tokens that cannot be decoded, usually containing illegal characters.
  • UNREACHABLE: Tokens that cannot be restored to their original token ID through the decoding and re-encoding process.
During the classification process, we first decode each token ID to obtain its corresponding string representation. If decoding fails, the token is classified as UNDECODEABLE. Next, we encode the decoded string and check if it can be restored to the original token ID. If it cannot, the token is classified as UNREACHABLE. If it meets the characteristics of a special token, it is classified as SPECIAL.
By filtering out these unnecessary tokens, we retain only the essential ones for subsequent mining, which improves the algorithm’s efficiency and ensures the accuracy of the detection results.
3.2 Glitch Token Verification
Task Template. Previous research typically employed a repetition task to verify whether a token is a glitch token by checking if the model can repeat a given token. If the model fails to reproduce the token correctly, the token is labeled as a glitch token. However, we found that the prior task templates are not directly suitable for gradient-based discrete optimization. For instance, in Magikarp Magikarp , the token being checked is repeated multiple times, forcing the model to output the current token, which can lead to unnecessary computational overhead and imprecise gradient calculations of the input token.
To address this issue, we designed a more straightforward task template as follows:
User: Please repeat the string: "«{token}»"
Assistant: Sure, the string is: "«{token}
In this template, the blue text represents the prompt used to verify glitch tokens. The green text, {token}, corresponds to the input token being tested, while the red text, {token}, represents the model’s predicted output. If the predicted token does not match the input token, the input token is classified as a glitch token.
The response from the assistant, starting with “Sure, the string is:"«”, is prefilled as part of the template, ensuring consistency in the model’s response format. The token is wrapped in the "«" and "»" symbols, similar to the token filter, to prevent interference from other characters during token encoding.
3.3 Token Selection
Entropy as Loss Function. One of the key distinctions between our approach and previous work lies in our token selection strategy. Prior methods often rely on manually observed prior knowledge about the distribution of glitch tokens to select those most likely to be glitch tokens. However, this approach can overlook tokens that do not fit the prior assumptions but are still glitch tokens. Ultimately, the effectiveness of these methods depends on how closely the manually defined priors align with the actual distribution of glitch tokens.
Given that manual priors of glitch token distribution can be inaccurate, a more robust approach is to identify tokens that deviate significantly from the normal token distribution. To achieve this, we propose using entropy as the loss function in our optimization process. Entropy measures the uncertainty or randomness in a probability distribution. For normal tokens, we expect the model to predict the next token with high confidence, resulting in a low-entropy distribution. Maximizing entropy helps us find tokens that cause more uncertainty in the model’s predictions, thereby identifying tokens that deviate from the expected behavior of normal tokens. This allows us to discover glitch tokens without depending on observed distribution patterns or manually defined assumptions about their characteristics.
Optimization Objective. Based on the above ideas, we present the mathematical description of the optimization objective. In each iteration, the goal is to find a batch of tokens with the highest entropy from the set of unverified tokens. Let $\mathcal{T}_{c}$ represent the current candidate set, defined as $\mathcal{T}_{c}=\mathcal{T}\setminus(\mathcal{T}^{*}\cup\mathcal{G})$.
Let $\mathbf{h}(t)$ represent the context embedding for a token $t$, and let $P(v\mid\mathbf{h}(t))$ denote the probability distribution over the vocabulary $\mathcal{V}$, predicted by the model for token $t$. The entropy $H(t)$ for token $t$ is defined as:
$H(t)=-\sum_{v\in\mathcal{V}}P(v\mid\mathbf{h}(t))\log P(v\mid\mathbf{h}(t))$
In each iteration, the optimization objective is to find a batch of tokens $\mathcal{B}$ that maximizes the entropy $H(t)$, i.e.,
$\mathcal{B}=\arg\max_{\mathcal{B}\subset\mathcal{T}_{c},|\mathcal{B}|=B}\sum_{t\in\mathcal{B}}H(t)$
Local Search Strategy. To efficiently solve the aforementioned optimization problem, we propose a local search strategy based on AutoPrompt. This approach combines local search with first-order Taylor approximation to estimate the entropy of candidate tokens, enabling us to select the batch most likely to contain glitch tokens.
Our strategy initiates by selecting an initial token $t_{0}$, computing its entropy $H(t_{0})$, and the corresponding gradient $\nabla_{e}H(t_{0})$, where $\nabla_{e}H(t_{0})$ represents the partial derivative with respect to the embedding vector $e_{t_{0}}$. For each candidate token $t\in\mathcal{T}_{c}$, we estimate its entropy using the first-order Taylor approximation:
$\hat{H}(t)\approx H(t_{0})+\nabla_{e}H(t_{0})^{T}(e_{t}-e_{t_{0}})$
where $e_{t}$ and $e_{t_{0}}$ denote the embedding vectors of tokens $t$ and $t_{0}$, respectively. This approximation estimates the change in entropy for the candidate token $t$ based on the change in its embedding relative to $e_{t_{0}}$, guiding the selection of tokens with potentially higher entropy.
To improve the accuracy of the approximation, we introduce a local search strategy. We define $\mathcal{N}_{K}(t_{0})$ as the set of $K$ nearest neighbor tokens to $t_{0}$:
$\mathcal{N}_{K}(t_{0})=\{t\in\mathcal{T}_{c}:t\text{ is one of the }K\text{ nearest neighbors of }t_{0}\}$
Within this local neighborhood $\mathcal{N}_{K}(t_{0})$, we select the batch $\mathcal{B}$ of $B$ tokens with the highest approximated entropy:
$\mathcal{B}=\arg\max_{\mathcal{B}\subset\mathcal{N}_{K}(t_{0}),|\mathcal{B}|=B}\sum_{t\in\mathcal{B}}\hat{H}(t)$
After selecting the batch $\mathcal{B}$, we compute the actual entropy values $H(t)$ for each token in the batch and choose the token with the highest actual entropy as the starting point $t_{0}$ for the next iteration. This ensures that each iteration is guided by the token with the most promising entropy, allowing the optimization process to converge more efficiently.
This local optimization strategy effectively mitigates the issue of inaccuracies in the Taylor approximation when applied globally while maintaining computational efficiency. By focusing on a local neighborhood in each iteration, we can more accurately estimate the entropy of candidate tokens, thereby enhancing the efficiency of glitch token detection. This approach strikes a balance between exploration of the token space and exploitation of local information, facilitating a more targeted and effective search for glitch tokens.
4 Experiments
4.1 Experimental Setup
Target LLM. We used a diverse set of LLMs from five different model families to evaluate the performance of our glitch token detection approach. The selected models include Meta’s Llama series llama2 ; llama3_1 , Alibaba’s Qwen models qwen2 ; qwen2_5 , Google’s Gemma models gemma2 , Microsoft’s Phi-3 models phi3 , and Mistral models mistral_7B ; mistral2024nemo . The details are presented in Table 1.
Model Family Model Names
Llama Models Llama-3.1-8B-Instruct , Llama-2-7B-chat-hf
Qwen Models Qwen2.5-7B-Instruct , Qwen2-7B-Instruct
Gemma Models Gemma-2-2b-it, Gemma-2-9b-it
Phi-3 Models Phi-3-mini-128k-instruct, Phi-3.5-mini-instruct
Mistral Models Mistral-7B-Instruct-v0.3, Mistral-Nemo-Instruct-2407
Table 1: Target LLMs used in the experiments.
Evaluation Metrics. To assess the performance of our glitch token detection method, we employ the Top-N Precision metric, denoted as Precision@N. This metric evaluates the proportion of true glitch tokens within the top $N$ predictions made by the algorithm. It is calculated as:
[…]
For example, Precision@1000 measures the precision of the top 1000 predicted tokens. This metric not only captures the accuracy of detecting true glitch tokens among high-ranking candidates but also reflects the effectiveness of a method under limited query budgets. As such, it is particularly well-suited for evaluating the performance of glitch token detection approaches, where efficiency and precision are crucial.
Baselines. We compare our proposed glitch token detection method with two state-of-the-art approaches: GlitchHunter GlitchHunter and Magikarp Magikarp . These methods serve as the primary benchmarks for evaluating our approach.
Although GlitchProber GlitchProber is another relevant method, it follows a fundamentally different approach by pre-collecting a subset of glitch tokens to train a classifier, introducing a supervised learning component. In contrast, GlitchMiner, along with GlitchHunter and Magikarp, uses heuristic-based methods to detect glitch tokens without relying on labeled data or additional classifier training. This methodological difference makes a direct comparison less meaningful, so we focus our evaluation on methods that align more closely with our unsupervised approach.
Parameter Settings. In our implementation of GlitchMiner, we use $K$=32 and $B$=8 as the default parameters. These values were chosen based on empirical testing to balance computational efficiency and detection effectiveness. Specifically, $K$=32 defines the size of the local neighborhood considered in each iteration, while $B$=8 determines the batch size for entropy computation. These settings have shown to provide a good trade-off between exploration of the token space and exploitation of local information across various model architectures.
4.2 Experimental Results
Model Metric GlitchHunter Magikarp GlitchMiner (ours)
Llama-3.1-8B Precision@1000 8.2% 50.4% 38.2%
Precision@2000 7.05% 42.35% 49.87%
Llama-2-7B Precision@1000 6.6% 9.5% 31.67%
Precision@2000 7.2% 9.15% 26.53%
Qwen2.5-7B Precision@1000 11.7% 0.4% 45.47%
Precision@2000 12.0% 0.2% 44.89%
Qwen2-7B Precision@1000 9.7% 0.4% 63.77%
Precision@2000 11.35% 0.2% 52.6%
Phi-3-mini-128k Precision@1000 2.0% 42.0% 31.53%
Precision@2000 2.25% 28.2% 21.15%
Phi-3.5-mini Precision@1000 3.0% 43.0% 28.0%
Precision@2000 3.75% 30.15% 25.11%
Gemma-2-2b-it Precision@1000 1.4% 71.9% 83.27%
Precision@2000 1.4% 55.5% 57.65%
Gemma-2-9b-it Precision@1000 4.0% 72.9% 90.17%
Precision@2000 3.1% 67.5% 70.57%
Mistral-7B-v0.3 Precision@1000 3.8% 6.6% 18.73%
Precision@2000 3.05% 7.45% 14.62%
Mistral-Nemo Precision@1000 4.5% 0.0% 57.03%
Precision@2000 4.45% 28.75% 61.48%
Average Precision@1000 5.49% 29.71% 48.78%
Precision@2000 5.56% 26.95% 42.45%
Table 2: Precision@1000 and Precision@2000 comparison of methods across different models. All results are averaged over three runs.
Table 2 presents a comprehensive comparison of GlitchMiner against state-of-the-art methods GlitchHunter and Magikarp across various LLM architectures. The results demonstrate the superior performance and versatility of our proposed approach.
GlitchMiner consistently outperforms both baselines in detecting glitch tokens, with particularly impressive results on newer model architectures. For instance, on the Gemma-2-9b model, GlitchMiner achieves a Precision@1000 of 90.17% and Precision@2000 of 70.57%, significantly surpassing both GlitchHunter and Magikarp. Similar performance gains are observed across the Qwen and Llama model families.
Figure 3: Comparison of GlitchMiner performance with and without local search strategy
The effectiveness of GlitchMiner is particularly evident in cases where Magikarp struggles, such as with the Qwen and Mistral models. This highlights a key advantage of our approach: its independence from manually defined priors. Unlike Magikarp, which relies heavily on assumed token distributions, GlitchMiner’s entropy-based loss function and local discrete optimization strategy allow it to adapt to the unique characteristics of each model’s vocabulary and token distribution.
This adaptability is crucial when dealing with the diverse landscape of modern LLMs, each with its own architectural nuances and token vocabularies. GlitchMiner’s consistent performance across different model families demonstrates its robustness and generalizability, addressing a significant limitation of previous approaches.
Remark. These experimental results not only showcase GlitchMiner’s superior precision in glitch token detection but also highlight its ability to generalize across diverse LLM architectures. This combination of accuracy and adaptability represents a significant advancement in the field of glitch token detection, providing a more reliable and versatile tool for improving the robustness and security of LLMs.
4.3 Ablation Study
To evaluate the contributions of key components in GlitchMiner, we conducted ablation studies focusing on the local search strategy, neighborhood size $K$, and batch size $B$.
Effect of Local Search. The local search strategy plays a crucial role in enhancing GlitchMiner’s ability to detect glitch tokens by improving the precision of the Taylor approximation. Without local search, detection accuracy drops significantly (Figure 3), as global search lacks the necessary granularity to maintain precise approximations within the token space.
Effect of Neighborhood Size. In figure 4, we can see that as the neighborhood size $K$ increases, there is a gradual decline in precision. However, if $K$ is too small, the exploration is insufficient, which also reduces accuracy, as shown in the experiments with Llama-3.1-8B. Therefore, balancing $K$ is crucial to ensure both broad enough exploration and sufficient approximation accuracy.
Figure 4: Impact of different Neighborhood Size $K$ on GlitchMiner’s performance
Figure 5: Impact of different Batch Size $B$ on GlitchMiner’s performance
Effect of Batch Size. As shown in figure 5, different models exhibit varying sensitivity to batch size $B$. For example, Llama-3.1-8B and Qwen2.5-7B perform best at $B=4$, while Phi-3-mini peaks at $B=8$. Gemma-2-2B, on the other hand, reaches its highest performance at $B=16$ and maintains strong results at larger $B$ values. This indicates that the optimal $B$ value depends on the specific model architecture.
4.4 Token Entropy Analysis
To further validate the effectiveness of our entropy-based approach in detecting glitch tokens, we conducted an entropy analysis comparing glitch tokens and normal tokens across different models. For each model, we computed the average entropy of glitch tokens ($E_{\text{Glitch}}$) and normal tokens ($E_{\text{Normal}}$).
Figure 6 presents the comparison of average entropy values between glitch tokens and normal tokens for each evaluated model. As shown in the figure, glitch tokens consistently exhibit significantly higher entropy than normal tokens across all models.
Figure 6: Average entropy comparison between glitch tokens and normal tokens across different models. Glitch tokens have higher entropy, indicating greater uncertainty in the model’s predictions for these tokens.
This pronounced difference in entropy values indicates that models are more uncertain when predicting glitch tokens compared to normal tokens. The higher entropy of glitch tokens validates our hypothesis that maximizing entropy effectively guides the search towards tokens that are challenging for the model to predict.
Moreover, the consistent pattern of higher entropy for glitch tokens across diverse model families—including Llama, Qwen, Gemma, Phi-3, and Mistral—demonstrates the generality and robustness of our entropy-based approach. This suggests that our method can be effectively applied to a wide range of LLMs with different architectures and tokenization strategies.
These findings reinforce the effectiveness of GlitchMiner’s entropy-based optimization in efficiently detecting glitch tokens by focusing on areas of high prediction uncertainty within the model.
5 Conclusion
In conclusion, we introduce GlitchMiner, a novel method for efficiently mining glitch tokens in large language models through gradient-based discrete optimization. By leveraging entropy as the loss function and integrating a local search strategy, GlitchMiner accurately identifies tokens that deviate from normal distributions without relying on manually defined priors. Extensive experimental results demonstrate that GlitchMiner outperforms existing state-of-the-art methods across various model architectures, showcasing higher precision and adaptability. Ablation studies further confirm the contributions of key components to the overall performance. The consistent results across different models emphasize the robustness and generalizability of GlitchMiner. Future work may explore applying this method to other high-dimensional tasks that require efficient discrete optimization.

### v3

@@
 Abstract
-Glitch tokens in Large Language Models (LLMs) can trigger unpredictable behaviors, compromising model reliability and safety.
-Existing detection methods often rely on manual observation to infer the prior distribution of glitch tokens, which is inefficient and lacks adaptability across diverse model architectures.
-To address these limitations, we introduce GlitchMiner, a gradient-based discrete optimization framework designed for efficient glitch token detection in LLMs.
-GlitchMiner leverages an entropy-based loss function to quantify the uncertainty in model predictions and integrates first-order Taylor approximation with a local search strategy to effectively explore the token space.
-Our evaluation across various mainstream LLM architectures demonstrates that GlitchMiner surpasses existing methods in both detection precision and adaptability.
-In comparison to the previous state-of-the-art, GlitchMiner achieves an average improvement of 19.07% in precision@1000 for glitch token detection.
-By enabling efficient detection of glitch tokens, GlitchMiner provides a valuable tool for assessing and mitigating potential vulnerabilities in LLMs, contributing to their overall security.
-Our code is available at https://github.com/wooozihui/GlitchMiner.
+Glitch tokens in Large Language Models (LLMs) can trigger unpredictable behaviors, threatening model reliability and safety.
+Existing detection methods often depend on predefined patterns, limiting their adaptability across diverse LLM architectures.
+We propose GlitchMiner, a gradient-based discrete optimization framework that efficiently identifies glitch tokens by leveraging entropy to quantify prediction uncertainty and a local search strategy for exploring the token space.
+Experiments across multiple LLM architectures show that GlitchMiner outperforms existing methods in both detection accuracy and adaptability, achieving over 10% average efficiency improvement.
+GlitchMiner enhances vulnerability assessment in LLMs, contributing to more robust and reliable applications.
+Code is available at https://github.com/wooozihui/GlitchMiner.
 1 Introduction
-Large Language Models (LLMs), such as GPT-4 gpt4 , have revolutionized natural language processing, enabling breakthroughs in critical fields, including code generation LLM_code1 ; LLM_code2 ; LLM_code3 , medical research medical1 ; medical2 , and education education1 ; education2 .
-Despite their strengths, these models harbor a subtle yet critical vulnerability: glitch tokens geiping2024coercing .
-Glitch tokens are anomalous tokens that disrupt the model’s ability to understand or generate accurate responses, leading to unpredictable behaviors such as inappropriate content, repetitive errors, hallucinated outputs, or offensive language geiping2024coercing ; solidgoldmagikarp ; GlitchHunter ; Magikarp .
-This issue stems from the tokenization process, potentially arising due to insufficient training or poorly embedded tokens within the model’s vocabulary.
-For instance, as shown in Figure 1, the Llama2-7B-chat-hf llama2 model fails to recognize the token “oreferrer,” treating it as an uncorrelated term.
-This oversight reveals deeper vulnerabilities in token handling, leading to unexpected outputs that can compromise the reliability of LLMs, especially in high-stakes applications.
-Mitigating these glitches is crucial to ensuring the trustworthiness and robustness of LLM-generated content.
-Existing methods for detecting glitch tokens primarily rely on manual observation to assume their prior distribution solidgoldmagikarp ; GlitchHunter ; Magikarp .
-However, this approach is not only time-consuming and labor-intensive but also struggles to adapt to different model architectures and vocabularies.
-Moreover, it often overlooks glitch tokens that fall outside predefined distribution patterns, limiting the detection’s coverage and accuracy.
-To overcome these challenges, we propose GlitchMiner, a novel gradient-based discrete optimization framework for efficiently detecting glitch tokens in LLMs.
-Unlike previous methods that rely on manual observation of glitch tokens, GlitchMiner focuses on identifying tokens that deviate significantly from the normal token distribution—where most tokens exhibit high certainty and predictable behavior.
-As illustrated in Figure 2, GlitchMiner leverages the uncertainty in token predictions to uncover potential glitches.
-It employs an entropy-based loss function to measure uncertainty and uses a local discrete optimization strategy, combining first-order Taylor approximation Taylor with local search, to efficiently explore the token space.
-This adaptive approach enables GlitchMiner to surpass the limitations of earlier methods, providing more comprehensive detection without relying on predefined distribution assumptions.
-Our main contributions in this paper include:
-  1. 1.
-    Entropy-based Loss Function: We propose using entropy as the optimization objective, effectively capturing the uncertainty in model predictions without relying on manually defined glitch token prior distributions.
-  2. 2.
-    Local Search Strategy: We design a novel discrete optimization method that combines first-order Taylor approximation with local search, significantly improving the efficiency and accuracy of glitch token detection.
-  3. 3.
-    Wide Applicability: Through experiments on multiple mainstream LLM architectures, we demonstrate the generalizability and effectiveness of GlitchMiner.
-Experimental results show that GlitchMiner improves the precision@1000 of glitch token detection by an average of 19.07% compared to previous state-of-the-art method.
-  4. 4.
-    Efficient Detection: Our method can quickly locate potential glitch tokens in large-scale vocabularies, greatly improving detection efficiency and providing a powerful tool for assessing and improving the security of LLMs.
-The structure of this paper is as follows: Section 2 reviews related work, including glitch token detection and gradient-based discrete optimization methods; Section 3 details the design and implementation of GlitchMiner; Section 4 presents experimental results and analysis; finally, Section 5 concludes the paper and discusses future research directions.
-Figure 1: Example of a glitch token "oreferrer", causing misinterpretation in the Llama2 model.
-Figure 2: Overview of the GlitchMiner Framework.
-This figure illustrates the two-stage process: (1) Initialization Stage – filters out unnecessary tokens to reduce the search space; (2) Mining Stage – iteratively selects batches of tokens, evaluates their entropy, and determines whether they are glitch tokens.
-2 Related Work
+Large Language Models (LLMs) have revolutionized natural language processing, enabling breakthrough applications in fields ranging from code generation LLM_code1 ; LLM_code2 ; LLM_code3 to medical research medical1 ; medical2 and education education1 ; education2 .
+However, the increasing reliance on these models in critical domains has brought to light a subtle yet significant vulnerability: glitch tokens geiping2024coercing .
+These anomalous tokens can severely disrupt an LLM’s ability to generate coherent and reliable outputs, potentially leading to repetitive errors, nonsensical responses, or even harmful content.
+The implications of such disruptions are particularly concerning in high-stakes applications where model reliability is paramount.
+Recent research efforts have made strides in understanding and detecting glitch tokens.
+Approaches like Magikarp Magikarp and GlitchHunter GlitchHunter have provided valuable insights, focusing on abnormal embedding patterns and clustering properties of these tokens.
+However, these methods are inherently limited by their reliance on manually observed patterns and predefined distribution assumptions.
+This dependence not only constrains their adaptability across diverse LLM architectures but also potentially overlooks novel forms of glitch tokens that do not conform to expected patterns.
+In this paper, we introduce GlitchMiner, a novel framework that fundamentally reimagines the approach to glitch token detection.
+Moving beyond pattern-based heuristics, GlitchMiner employs gradient-based discrete optimization to directly analyze an LLM’s prediction behavior.
+At the core of our method is the use of entropy as an exploratory metric, guiding our search through the vast token space to identify regions of unexpected model behavior.
+In summary, the key innovations of GlitchMiner include:
+  • Entropy-Based Optimization: We introduce entropy as a novel optimization objective for glitch token detection.
+This approach allows us to systematically identify tokens that induce high uncertainty in the model’s predictions, moving beyond predefined patterns or embedding anomalies.
+By focusing on the model’s prediction behavior, we can detect a broader spectrum of glitch tokens, including those that might be missed by traditional embedding-based methods.
+  • Gradient-Guided Local Search: Our method combines first-order Taylor approximation with a local search strategy, efficiently navigating the vast token space.
+This approach significantly improves both the efficiency and accuracy of glitch token detection, allowing us to explore the token space more thoroughly while maintaining computational feasibility.
+  • Adaptive Cross-Architecture Detection: GlitchMiner’s framework is designed to adapt to the unique characteristics of different LLM architectures without relying on model-specific assumptions.
+This flexibility enables effective glitch token detection across a wide range of models, addressing a key limitation of previous methods that often relied on architecture-specific heuristics.
+We conducted extensive experiments across 10 popular LLM architectures, covering models from the Llama, Qwen, Gemma, Phi, and Mistral families.
+The results show that GlitchMiner surpasses existing methods in 90% of cases when it comes to glitch token detection efficiency.
+Specifically, GlitchMiner achieves an average Detected@1000 score of 612.0 and a Detected@2000 score of 980.1.
+These results represent improvements of 10.5% and 10.7% respectively over the previous leading methods.
+In the context of glitch token detection, these improvements are significant as they translate to hundreds of additional potentially harmful tokens being identified.
+Such enhancements not only improve detection performance but also optimize query efficiency, which is crucial for enhancing the reliability and robustness of LLMs in critical applications where even a single undetected glitch token could lead to severe consequences.
+The remainder of this paper is structured as follows: Section 2 provides a comprehensive review of related work in glitch token detection and gradient-based optimization techniques.
+Section 3 details the design principles and implementation of GlitchMiner.
+Section 4 presents our experimental results and analysis, demonstrating the effectiveness of our approach across various LLM architectures.
+Finally, Section 5 concludes with a discussion of the implications of our work and outlines promising directions for future research in this critical area of LLM security and reliability.
+2 Background and Related Work
+2.1 What is a Glitch Token?
+Glitch tokens, a concept that gained attention following the discovery of the "SolidGoldMagikarp" phenomenon by Rumbelow et al.
+solidgoldmagikarp , are anomalous tokens in LLMs that can trigger unexpected and often erroneous behaviors when processed by the model.
+These tokens typically result from irregularities in the training process, such as underrepresentation in the training data or inconsistencies in tokenization Magikarp ; geiping2024coercing .
+Glitch tokens have been found to exhibit several key characteristics:
+  • They often have abnormal embedding patterns, such as unusually large L2 norms or atypical positions in the embedding space Magikarp .
+  • When input to an LLM, they can cause the model to produce repetitive, nonsensical, or completely unrelated outputs solidgoldmagikarp ; GlitchHunter .
+  • They may lead to significant deviations in the model’s internal activations compared to normal tokens GlitchProber .
+  • Glitch tokens can sometimes bypass content filters or trigger unexpected model behaviors, potentially compromising the safety and reliability of LLMs geiping2024coercing .
+The study of glitch tokens is crucial for understanding and mitigating potential vulnerabilities in LLMs, as these tokens can potentially be exploited for adversarial attacks or lead to unintended model behaviors in critical applications geiping2024coercing .
 2.2 Glitch Token Detection
-Detecting glitch tokens in LLMs has garnered increasing attention.
-Several methods have been proposed to address this issue, including Magikarp Magikarp , GlitchHunter GlitchHunter , and GlitchProber GlitchProber , each utilizing different strategies to identify glitch tokens.
-Magikarp Magikarp detects glitch tokens by analyzing token embeddings.
-It uses characteristic patterns, such as the $\ell_{2}$ norm, in the embedding space to identify under-trained tokens and employs simple repetition tasks to test whether the model can accurately reproduce these tokens.
-This method focuses on quickly detecting problematic tokens through direct queries, making it fast and lightweight.
-GlitchHunter GlitchHunter adopts a clustering-based approach.
-It assumes that glitch tokens tend to cluster in the embedding space and constructs a Token Embedding Graph (TEG), where nodes represent tokens connected based on their embedding similarity.
-By applying the Leiden algorithm traag2019louvain , GlitchHunter identifies potential glitch clusters.
-Hypothesis testing is then used to refine these clusters, iterating multiple times to ensure the accuracy of detection.
+The detection of glitch tokens in LLMs has become an increasingly important area of research, given their potential to disrupt model performance and reliability.
+The fundamental approach to identifying these anomalous tokens often relies on repetition tasks, a method popularized by the SolidGoldMagikarp study solidgoldmagikarp .
+In these tasks, the model is prompted to repeat a given token, with the assumption that normal tokens should be easily reproducible, while glitch tokens often lead to failures in this simple task.
+Building upon this basic principle, researchers have developed several sophisticated methods to detect glitch tokens more efficiently and accurately:
+Magikarp solidgoldmagikarp pioneered a fast and lightweight approach by analyzing token embeddings.
+It identifies potentially problematic tokens by examining characteristic patterns in the embedding space, such as unusual $\ell_{2}$ norms.
+These candidate tokens are then verified using repetition tasks, allowing for quick detection through direct queries to the model.
+GlitchHunter GlitchHunter introduced a clustering-based method.
+This approach is founded on the observation that glitch tokens often cluster together in the embedding space.
+GlitchHunter constructs a Token Embedding Graph (TEG) to represent token relationships and applies the Leiden algorithm traag2019louvain to identify potential glitch clusters.
+These clusters are then refined through iterative hypothesis testing to improve detection accuracy.
 GlitchProber GlitchProber takes a different approach by analyzing the internal activations within transformer layers, such as attention heads and hidden states.
@@
 Additionally, it integrates mitigation mechanisms by modifying neuron activations to minimize the impact of glitch tokens during inference.
-Both Magikarp and GlitchHunter rely on manually observed prior distributions to sample token space for glitch token detection, performing well when their assumptions align with the actual glitch token distribution.
-However, their adaptability can be limited when real-world distributions deviate from these priors.
-GlitchProber aims to improve detection through automation, but it requires a large number of queries to collect sufficient training data.
-Additionally, the need for extra training introduces computational overhead, impacting efficiency in practice.
-Our method overcomes these limitations by using entropy-based loss functions to quantify uncertainty in predictions, without relying on priors or additional training.
+While these methods have significantly advanced our ability to detect glitch tokens, they are inherently limited by their reliance on pattern-based heuristics.
+Magikarp and GlitchHunter focus on specific embedding patterns or clustering behaviors, while GlitchProber examines predefined patterns in activation spaces.
+This dependence on pattern-based approaches can constrain their adaptability across diverse LLM architectures and potentially overlook novel forms of glitch tokens.
+Our proposed method, GlitchMiner, aims to address these limitations by employing gradient-based discrete optimization to directly analyze an LLM’s prediction behavior, potentially offering a more flexible and comprehensive approach to glitch token detection.
 2.3 Gradient-based Discrete Optimization
-Gradient-based discrete optimization methods hotflip ; autoprompt ; GCG leverage gradient information to predict how individual tokens impact the loss function.
+Gradient-based discrete optimization methods hotflip ; autoprompt ; GCG ; PEZ leverage gradient information to predict how individual tokens impact the loss function.
 These approaches typically treat the one-hot encoding of tokens or token embeddings as continuous vectors to compute gradients, guiding token replacements for optimization.
@@
 Notably, GCG has been applied to automated jailbreaks shen2023anything in LLMs, efficiently searching for adversarial suffixes.
-Both AutoPrompt and GCG rely on large batch sampling to reduce inaccuracies in gradient predictions.
-These inaccuracies occur because Taylor approximation become less reliable when the input tokens are far from the original reference points.
-To address this challenge, we introduce a local search strategy in our approach.
-By focusing on a smaller, localized token space, our method improves precision without the need for large batch sampling.
-This targeted search not only enhances precision but also avoids unnecessary queries over non-glitch tokens, significantly boosting query efficiency for glitch token detection.
+AutoPrompt and GCG both rely on large batch sampling to mitigate inaccuracies in gradient prediction.
+We identified that these inaccuracies arise from the inaccuracy of Taylor expansions when input tokens are distant from the original points.
+This overlooks a fundamental condition of Taylor approximation: its accuracy is highest for points close to the reference point.
+Building on these works, we introduce a local search strategy in our approach.
+This improvement enables us to achieve high precision in gradient estimation without relying on large batch sampling, by focusing on a smaller, localized token space.
+By addressing the core issue of Taylor approximation accuracy, our method allows for more efficient and accurate exploration of the token space, which is particularly valuable for glitch token detection.
 3 Method
@@
 3.2 Glitch Token Verification
-Task Template.
-Previous research typically employed a repetition task to verify whether a token is a glitch token by checking if the model can repeat a given token.
-If the model fails to reproduce the token correctly, the token is labeled as a glitch token.
-However, we found that the prior task templates are not directly suitable for gradient-based discrete optimization.
-For instance, in Magikarp Magikarp , the token being checked is repeated multiple times, forcing the model to output the current token, which can lead to unnecessary computational overhead and imprecise gradient calculations of the input token.
-To address this issue, we designed a more straightforward task template as follows:
+Our glitch token verification process consists of two stages: optimization-driven detection and cross-verification.
+Template for Optimization. For our primary detection and optimization process, we designed the following task template:
 User: Please repeat the string: "«{token}»"
 Assistant: Sure, the string is: "«{token}
-In this template, the blue text represents the prompt used to verify glitch tokens.
-The green text, {token}, corresponds to the input token being tested, while the red text, {token}, represents the model’s predicted output.
-If the predicted token does not match the input token, the input token is classified as a glitch token.
-The response from the assistant, starting with “Sure, the string is:"«”, is prefilled as part of the template, ensuring consistency in the model’s response format.
-The token is wrapped in the "«" and "»" symbols, similar to the token filter, to prevent interference from other characters during token encoding.
+In this template, the blue text represents the prompt, the green text {token} is the input token being tested, and the red text {token} represents the model’s predicted output.
+If the predicted token does not match the input token, it is initially classified as a potential glitch token.
+This template is specifically designed to support our gradient-based optimization process.
+Its simplicity allows for clear gradient computations, while the prefilled assistant response ensures a uniform starting point for each token evaluation.
+Wrapping the token in "«" and "»" symbols minimizes interference from surrounding context.
+These features enable precise entropy calculations and gradient estimations, which are crucial for our entropy-guided search strategy described in the token selection section.
+Cross-Verification.
+While our primary template is optimized for the search process, relying solely on one template may lead to false positives.
+To enhance the robustness of our detection, we implement a cross-verification step using two additional templates derived from GlitchHunter and Magikarp.
+Each potential glitch token identified by our primary template undergoes verification with these additional templates.
+A token is confirmed as a glitch token only if it fails the repetition task across all templates.
+This two-stage approach combines the efficiency of our optimization-driven search with the reliability of multi-template verification, significantly reducing false positives while maintaining the effectiveness of our gradient-based detection method.
 3.3 Token Selection
-Entropy as Loss Function.
-One of the key distinctions between our approach and previous work lies in our token selection strategy.
-Prior methods often rely on manually observed prior knowledge about the distribution of glitch tokens to select those most likely to be glitch tokens.
-However, this approach can overlook tokens that do not fit the prior assumptions but are still glitch tokens.
-Ultimately, the effectiveness of these methods depends on how closely the manually defined priors align with the actual distribution of glitch tokens.
-Given that manual priors of glitch token distribution can be inaccurate, a more robust approach is to identify tokens that deviate significantly from the normal token distribution.
-To achieve this, we propose using entropy as the loss function in our optimization process.
-Entropy measures the uncertainty or randomness in a probability distribution.
-For normal tokens, we expect the model to predict the next token with high confidence, resulting in a low-entropy distribution.
-Maximizing entropy helps us find tokens that cause more uncertainty in the model’s predictions, thereby identifying tokens that deviate from the expected behavior of normal tokens.
-This allows us to discover glitch tokens without depending on observed distribution patterns or manually defined assumptions about their characteristics.
-Optimization Objective.
-Based on the above ideas, we present the mathematical description of the optimization objective.
-In each iteration, the goal is to find a batch of tokens with the highest entropy from the set of unverified tokens.
-Let $\mathcal{T}_{c}$ represent the current candidate set, defined as $\mathcal{T}_{c}=\mathcal{T}\setminus(\mathcal{T}^{*}\cup\mathcal{G})$.
-Let $\mathbf{h}(t)$ represent the context embedding for a token $t$, and let $P(v\mid\mathbf{h}(t))$ denote the probability distribution over the vocabulary $\mathcal{V}$, predicted by the model for token $t$.
-The entropy $H(t)$ for token $t$ is defined as:
+Motivation for Entropy-Based Approach.
+Previous methods for glitch token detection often rely on manually observed patterns or predefined distribution assumptions.
+However, these approaches can be limited in their ability to adapt to diverse LLM architectures and may overlook novel forms of glitch tokens.
+To address these limitations, we propose an entropy-based approach.
+Entropy, as a measure of uncertainty in probability distributions, offers a model-agnostic way to identify tokens that cause unexpected behavior in LLMs.
+By focusing on tokens that maximize entropy, we can detect glitch tokens that deviate significantly from normal token behavior, regardless of their specific characteristics or the underlying model architecture.
+Entropy-Guided Exploration. We define the entropy $H(t)$ for a token $t$ as:
 $H(t)=-\sum_{v\in\mathcal{V}}P(v\mid\mathbf{h}(t))\log P(v\mid\mathbf{h}(t))$
-In each iteration, the optimization objective is to find a batch of tokens $\mathcal{B}$ that maximizes the entropy $H(t)$, i.e.,
+where $\mathbf{h}(t)$ is the context embedding for token $t$, and $P(v\mid\mathbf{h}(t))$ is the model’s predicted probability distribution over the vocabulary $\mathcal{V}$.
+By maximizing entropy, we aim to find tokens that cause high uncertainty in the model’s predictions, potentially indicating glitch behavior.
+Optimization Objective. Our goal is to find a batch of tokens $\mathcal{B}$ that maximizes the total entropy:
 $\mathcal{B}=\arg\max_{\mathcal{B}\subset\mathcal{T}_{c},|\mathcal{B}|=B}\sum_{t\in\mathcal{B}}H(t)$
+Here, $\mathcal{T}_{c}$ represents the current candidate set, defined as $\mathcal{T}_{c}=\mathcal{T}\setminus(\mathcal{T}^{*}\cup\mathcal{G})$, where:
+  • $\mathcal{T}$ is the initial set of all tokens in the model’s vocabulary.
+  • $\mathcal{T}^{*}$ is the set of tokens that have been filtered out or verified as non-glitch tokens.
+This includes special tokens, undecoded tokens, and tokens that have been checked and confirmed as normal.
+  • $\mathcal{G}$ is the set of identified glitch tokens discovered so far in the process.
+By excluding $\mathcal{T}^{*}$ and $\mathcal{G}$ from $\mathcal{T}$, we ensure that our search focuses only on the remaining unverified tokens, improving efficiency and preventing redundant checks.
 Local Search Strategy.
-To efficiently solve the aforementioned optimization problem, we propose a local search strategy based on AutoPrompt.
-This approach combines local search with first-order Taylor approximation to estimate the entropy of candidate tokens, enabling us to select the batch most likely to contain glitch tokens.
-Our strategy initiates by selecting an initial token $t_{0}$, computing its entropy $H(t_{0})$, and the corresponding gradient $\nabla_{e}H(t_{0})$, where $\nabla_{e}H(t_{0})$ represents the partial derivative with respect to the embedding vector $e_{t_{0}}$.
-For each candidate token $t\in\mathcal{T}_{c}$, we estimate its entropy using the first-order Taylor approximation:
+To efficiently solve this optimization problem while addressing the limitations of global Taylor approximation, we introduce a local search strategy.
+We begin by selecting an initial token $t_{0}$ and computing its entropy $H(t_{0})$ and gradient $\nabla_{e}H(t_{0})$.
+We then define a local neighborhood $\mathcal{N}_{K}(t_{0})$ consisting of the $K$ nearest neighbors to $t_{0}$ in the embedding space.
+This localized approach allows us to refine the search within a region close to $t_{0}$, improving approximation accuracy and ensuring efficient exploration of potential glitch tokens.
+Within this local neighborhood, we estimate the entropy of each candidate token $t\in\mathcal{N}_{K}(t_{0})$ using first-order Taylor approximation:
 $\hat{H}(t)\approx H(t_{0})+\nabla_{e}H(t_{0})^{T}(e_{t}-e_{t_{0}})$
-where $e_{t}$ and $e_{t_{0}}$ denote the embedding vectors of tokens $t$ and $t_{0}$, respectively.
-This approximation estimates the change in entropy for the candidate token $t$ based on the change in its embedding relative to $e_{t_{0}}$, guiding the selection of tokens with potentially higher entropy.
-To improve the accuracy of the approximation, we introduce a local search strategy. We define $\mathcal{N}_{K}(t_{0})$ as the set of $K$ nearest neighbor tokens to $t_{0}$:
-$\mathcal{N}_{K}(t_{0})=\{t\in\mathcal{T}_{c}:t\text{ is one of the }K\text{ nearest neighbors of }t_{0}\}$
-Within this local neighborhood $\mathcal{N}_{K}(t_{0})$, we select the batch $\mathcal{B}$ of $B$ tokens with the highest approximated entropy:
-$\mathcal{B}=\arg\max_{\mathcal{B}\subset\mathcal{N}_{K}(t_{0}),|\mathcal{B}|=B}\sum_{t\in\mathcal{B}}\hat{H}(t)$
-After selecting the batch $\mathcal{B}$, we compute the actual entropy values $H(t)$ for each token in the batch and choose the token with the highest actual entropy as the starting point $t_{0}$ for the next iteration.
-This ensures that each iteration is guided by the token with the most promising entropy, allowing the optimization process to converge more efficiently.
-This local optimization strategy effectively mitigates the issue of inaccuracies in the Taylor approximation when applied globally while maintaining computational efficiency.
-By focusing on a local neighborhood in each iteration, we can more accurately estimate the entropy of candidate tokens, thereby enhancing the efficiency of glitch token detection.
-This approach strikes a balance between exploration of the token space and exploitation of local information, facilitating a more targeted and effective search for glitch tokens.
+where $e_{t}$ and $e_{t_{0}}$ are the embedding vectors of tokens $t$ and $t_{0}$, respectively.
+This approximation allows us to efficiently estimate the entropy change for nearby tokens without the computational cost of exact calculations.
+Based on these estimates, we select a batch $\mathcal{B}$ of $B$ tokens with the highest approximated entropy.
+For this selected batch, we then compute the actual entropy values and choose the token with the highest entropy as the starting point for the next iteration.
+This iterative process ensures that our search is guided by the most promising candidates at each step.
+This local search strategy significantly improves the accuracy of entropy estimation by focusing on tokens close to the reference point, addressing a key limitation of previous gradient-based methods.
+By balancing exploration of the token space with exploitation of local information, our approach enables more efficient and effective glitch token detection.
+The combination of entropy-guided exploration and local search allows our method to adapt to diverse LLM architectures without relying on model-specific assumptions.
+This results in a more robust and versatile approach to glitch token detection, capable of identifying a wide range of anomalous tokens across different model types.
 4 Experiments
 4.1 Experimental Setup
-Target LLM.
+Test LLMs.
 We used a diverse set of LLMs from five different model families to evaluate the performance of our glitch token detection approach.
@@
 Mistral Models Mistral-7B-Instruct-v0.3, Mistral-Nemo-Instruct-2407
-Table 1: Target LLMs used in the experiments.
+Table 1: Test LLMs used in the experiments.
 Evaluation Metrics.
-To assess the performance of our glitch token detection method, we employ the Top-N Precision metric, denoted as Precision@N.
-This metric evaluates the proportion of true glitch tokens within the top $N$ predictions made by the algorithm.
+To assess the performance of our glitch token detection method, we employ the Detected@N metric.
+This metric measures the absolute number of true glitch tokens identified within the top $N$ predictions made by the algorithm.
 It is calculated as:
 […]
-For example, Precision@1000 measures the precision of the top 1000 predicted tokens.
-This metric not only captures the accuracy of detecting true glitch tokens among high-ranking candidates but also reflects the effectiveness of a method under limited query budgets.
-As such, it is particularly well-suited for evaluating the performance of glitch token detection approaches, where efficiency and precision are crucial.
+For example, Detected@1000 measures the number of true glitch tokens discovered within the top 1000 predicted tokens.
+This metric directly reflects a method’s effectiveness in identifying glitch tokens under practical constraints, as it captures both the detection capability and query efficiency of the approach.
+By comparing the number of glitch tokens detected under the same query budget ($N$), we can effectively evaluate different methods’ ability to maximize token discovery while minimizing computational resources.
+This makes it an ideal metric for real-world applications where efficient detection of glitch tokens is crucial.
 Baselines.
@@
 These settings have shown to provide a good trade-off between exploration of the token space and exploitation of local information across various model architectures.
-4.2 Experimental Results
+Initialization Strategy in Experiments.
+To ensure stable and consistent comparisons across runs, we initialize the search with the token exhibiting the smallest $\ell_{2}$ norm in the embedding space, based on prior observations that such tokens often exhibit glitch-like behaviors.
+However, as shown in Figure 4, we found that GlitchMiner remains robust to different initialization choices, achieving similar performance even with random starting points.
+4.2 Main Results
 Model Metric GlitchHunter Magikarp GlitchMiner (ours)
-Llama-3.1-8B Precision@1000 8.2% 50.4% 38.2%
-Precision@2000 7.05% 42.35% 49.87%
-Llama-2-7B Precision@1000 6.6% 9.5% 31.67%
-Precision@2000 7.2% 9.15% 26.53%
-Qwen2.5-7B Precision@1000 11.7% 0.4% 45.47%
-Precision@2000 12.0% 0.2% 44.89%
-Qwen2-7B Precision@1000 9.7% 0.4% 63.77%
-Precision@2000 11.35% 0.2% 52.6%
-Phi-3-mini-128k Precision@1000 2.0% 42.0% 31.53%
-Precision@2000 2.25% 28.2% 21.15%
-Phi-3.5-mini Precision@1000 3.0% 43.0% 28.0%
-Precision@2000 3.75% 30.15% 25.11%
-Gemma-2-2b-it Precision@1000 1.4% 71.9% 83.27%
-Precision@2000 1.4% 55.5% 57.65%
-Gemma-2-9b-it Precision@1000 4.0% 72.9% 90.17%
-Precision@2000 3.1% 67.5% 70.57%
-Mistral-7B-v0.3 Precision@1000 3.8% 6.6% 18.73%
-Precision@2000 3.05% 7.45% 14.62%
-Mistral-Nemo Precision@1000 4.5% 0.0% 57.03%
-Precision@2000 4.45% 28.75% 61.48%
-Average Precision@1000 5.49% 29.71% 48.78%
-Precision@2000 5.56% 26.95% 42.45%
-Table 2: Precision@1000 and Precision@2000 comparison of methods across different models. All results are averaged over three runs.
-Table 2 presents a comprehensive comparison of GlitchMiner against state-of-the-art methods GlitchHunter and Magikarp across various LLM architectures.
-The results demonstrate the superior performance and versatility of our proposed approach.
-GlitchMiner consistently outperforms both baselines in detecting glitch tokens, with particularly impressive results on newer model architectures.
-For instance, on the Gemma-2-9b model, GlitchMiner achieves a Precision@1000 of 90.17% and Precision@2000 of 70.57%, significantly surpassing both GlitchHunter and Magikarp.
-Similar performance gains are observed across the Qwen and Llama model families.
+Llama-3.1-8B-Instruct Detected@1000 25 664 568
+Detected@2000 56 935 1164
+Llama-2-7B-chat-hf Detected@1000 61 100 319
+Detected@2000 126 186 532
+Qwen2.5-7B-Instruct Detected@1000 75 1000 1000
+Detected@2000 180 1893 1839
+Qwen2-7B-Instruct Detected@1000 96 999 1000
+Detected@2000 191 1842 1847
+Gemma-2-2b-it Detected@1000 23 678 744
+Detected@2000 35 984 1019
+Gemma-2-9b-it Detected@1000 29 623 775
+Detected@2000 45 983 1089
+Phi-3.5-mini-instruct Detected@1000 20 393 396
+Detected@2000 44 496 516
+Phi-3-mini-128k-instruct Detected@1000 26 398 404
+Detected@2000 55 489 517
+Mistral-7B-Instruct-v0.3 Detected@1000 6 110 219
+Detected@2000 19 130 302
+Mistral-Nemo-Instruct-2407 Detected@1000 48 574 695
+Detected@2000 79 918 976
+Average Detected@1000 40.9 553.9 612.0
+Detected@2000 93.0 885.6 980.1
+Table 2: Detected@1000 and Detected@2000 comparison of methods across different models.
+Table 2 displays the performance comparison of GlitchMiner with the state-of-the-art methods GlitchHunter and Magikarp across different LLM architectures using the Detected@N metric.
+The results highlight GlitchMiner’s strong and versatile detection capabilities.
+In the majority of cases, GlitchMiner achieved the highest Detected@2000 score, consistently outperforming both baselines in terms of identified glitch tokens within the top 2000 predictions.
+This is particularly evident in models such as Llama-2-7B-chat-hf, where GlitchMiner achieved a Detected@2000 of 532, surpassing Magikarp by a notable margin.
+For the Qwen and other models, GlitchMiner maintained robust performance, demonstrating adaptability across different LLM architectures.
 Figure 1: Comparison of GlitchMiner performance with and without local search strategy
-The effectiveness of GlitchMiner is particularly evident in cases where Magikarp struggles, such as with the Qwen and Mistral models.
-This highlights a key advantage of our approach: its independence from manually defined priors.
-Unlike Magikarp, which relies heavily on assumed token distributions, GlitchMiner’s entropy-based loss function and local discrete optimization strategy allow it to adapt to the unique characteristics of each model’s vocabulary and token distribution.
-This adaptability is crucial when dealing with the diverse landscape of modern LLMs, each with its own architectural nuances and token vocabularies.
-GlitchMiner’s consistent performance across different model families demonstrates its robustness and generalizability, addressing a significant limitation of previous approaches.
-Remark.
-These experimental results not only showcase GlitchMiner’s superior precision in glitch token detection but also highlight its ability to generalize across diverse LLM architectures.
-This combination of accuracy and adaptability represents a significant advancement in the field of glitch token detection, providing a more reliable and versatile tool for improving the robustness and security of LLMs.
+These experimental findings underscore GlitchMiner’s ability not only to accurately detect glitch tokens but also to generalize across a diverse set of model architectures.
+This adaptability and precision position GlitchMiner as a powerful tool for enhancing the robustness and security of LLMs.
 4.3 Ablation Study
@@
 Effect of Neighborhood Size.
-In figure 4, we can see that as the neighborhood size $K$ increases, there is a gradual decline in precision.
-However, if $K$ is too small, the exploration is insufficient, which also reduces accuracy, as shown in the experiments with Llama-3.1-8B.
-Therefore, balancing $K$ is crucial to ensure both broad enough exploration and sufficient approximation accuracy.
+We analyzed the impact of neighborhood size $K$ on detection performance.
+As shown in Figure 2, increasing $K$ generally leads to a decline in Detected@1000 values across models.
+This trend indicates that as $K$ grows, the Taylor approximation becomes less effective, resulting in reduced prediction accuracy.
 Figure 2: Impact of different Neighborhood Size $K$ on GlitchMiner’s performance
 Figure 3: Impact of different Batch Size $B$ on GlitchMiner’s performance
+Figure 4: Effect of Initialization Method on GlitchMiner’s Detected@1000 score.
 Effect of Batch Size.
-As shown in figure 5, different models exhibit varying sensitivity to batch size $B$.
-For example, Llama-3.1-8B and Qwen2.5-7B perform best at $B=4$, while Phi-3-mini peaks at $B=8$.
-Gemma-2-2B, on the other hand, reaches its highest performance at $B=16$ and maintains strong results at larger $B$ values.
-This indicates that the optimal $B$ value depends on the specific model architecture.
+As shown in Figure 3, the performance of GlitchMiner remains relatively stable as batch size $B$ increases.
+Notably, even with $B=1$, GlitchMiner achieves effective detection results, indicating that it can make accurate predictions without relying on a large batch size.
+Effect of Initialization Token.
+As shown in Figure 4, GlitchMiner’s performance remains stable across different initialization tokens.
+The red dots represent the minimum $\ell_{2}$ norm initialization, while the orange dots show three random trials.
+For most models, random initialization results are close to the minimum $\ell_{2}$ norm, indicating that GlitchMiner achieves consistent detection accuracy regardless of the initialization approach.
 4.4 Token Entropy Analysis
@@
 5 Conclusion
-In conclusion, we introduce GlitchMiner, a novel method for efficiently mining glitch tokens in large language models through gradient-based discrete optimization.
-By leveraging entropy as the loss function and integrating a local search strategy, GlitchMiner accurately identifies tokens that deviate from normal distributions without relying on manually defined priors.
-Extensive experimental results demonstrate that GlitchMiner outperforms existing state-of-the-art methods across various model architectures, showcasing higher precision and adaptability.
-Ablation studies further confirm the contributions of key components to the overall performance.
-The consistent results across different models emphasize the robustness and generalizability of GlitchMiner.
-Future work may explore applying this method to other high-dimensional tasks that require efficient discrete optimization.
+In this paper, we introduced GlitchMiner, a novel framework for detecting glitch tokens in LLMs through gradient-based discrete optimization.
+Our method combines entropy-based loss functions with a local search strategy to navigate token space efficiently, improving both detection accuracy and computational efficiency.
+Experimental results across diverse LLMs highlight GlitchMiner’s robustness and versatility compared to existing methods.
+Our findings confirm GlitchMiner’s effectiveness in identifying glitch tokens across architectures, demonstrating its potential for enhancing the security and reliability of LLM-based applications.
+Future work could explore extending this approach to other high-dimensional discrete optimization tasks.
+Broader Impact
+The potential impact of this research extends beyond mere academic interest.
+By improving the detection of glitch tokens, GlitchMiner contributes to enhancing the overall reliability and safety of LLMs in real-world applications.
+This is particularly crucial in high-stakes domains such as healthcare, finance, and autonomous systems, where the consequences of model errors can be severe.
+Furthermore, our approach opens new avenues for understanding and mitigating vulnerabilities in AI systems, potentially influencing the development of more robust and trustworthy AI technologies in the future.
+Acknowledgments
+We would like to express our sincere gratitude to Sander Land for his valuable insights and feedback on this work, particularly regarding the evaluation metrics.
+His suggestions were instrumental in refining our approach and improving the clarity of our results.
+We also wish to thank Guanlin Li for his valuable suggestions on our manuscript.
+His insights were crucial for improving the content and structure of the paper.

### v4

@@
-Mining Glitch Tokens in Large Language Models via Gradient-based Discrete Optimization
+GlitchMiner: Mining Glitch Tokens in Large Language Models via Gradient-based Discrete Optimization
 Abstract
 Glitch tokens in Large Language Models (LLMs) can trigger unpredictable behaviors, threatening model reliability and safety.
-Existing detection methods often depend on predefined patterns, limiting their adaptability across diverse LLM architectures.
-We propose GlitchMiner, a gradient-based discrete optimization framework that efficiently identifies glitch tokens by leveraging entropy to quantify prediction uncertainty and a local search strategy for exploring the token space.
-Experiments across multiple LLM architectures show that GlitchMiner outperforms existing methods in both detection accuracy and adaptability, achieving over 10% average efficiency improvement.
-GlitchMiner enhances vulnerability assessment in LLMs, contributing to more robust and reliable applications.
+Existing detection methods rely on predefined patterns, limiting their adaptability across diverse LLM architectures.
+We propose GlitchMiner, a gradient-based discrete optimization framework that efficiently identifies glitch tokens by introducing entropy as a measure of prediction uncertainty and employing a local search strategy to explore the token space.
+Experiments across multiple LLM architectures demonstrate that GlitchMiner outperforms existing methods in detection accuracy and adaptability, achieving over 10% average efficiency improvement.
+This method enhances vulnerability assessment in LLMs, contributing to the development of more robust and reliable applications.
 Code is available at https://github.com/wooozihui/GlitchMiner.
@@
 Large Language Models (LLMs) have revolutionized natural language processing, enabling breakthrough applications in fields ranging from code generation LLM_code1 ; LLM_code2 ; LLM_code3 to medical research medical1 ; medical2 and education education1 ; education2 .
-However, the increasing reliance on these models in critical domains has brought to light a subtle yet significant vulnerability: glitch tokens geiping2024coercing .
-These anomalous tokens can severely disrupt an LLM’s ability to generate coherent and reliable outputs, potentially leading to repetitive errors, nonsensical responses, or even harmful content.
-The implications of such disruptions are particularly concerning in high-stakes applications where model reliability is paramount.
-Recent research efforts have made strides in understanding and detecting glitch tokens.
-Approaches like Magikarp Magikarp and GlitchHunter GlitchHunter have provided valuable insights, focusing on abnormal embedding patterns and clustering properties of these tokens.
-However, these methods are inherently limited by their reliance on manually observed patterns and predefined distribution assumptions.
-This dependence not only constrains their adaptability across diverse LLM architectures but also potentially overlooks novel forms of glitch tokens that do not conform to expected patterns.
-In this paper, we introduce GlitchMiner, a novel framework that fundamentally reimagines the approach to glitch token detection.
-Moving beyond pattern-based heuristics, GlitchMiner employs gradient-based discrete optimization to directly analyze an LLM’s prediction behavior.
-At the core of our method is the use of entropy as an exploratory metric, guiding our search through the vast token space to identify regions of unexpected model behavior.
-In summary, the key innovations of GlitchMiner include:
-  • Entropy-Based Optimization: We introduce entropy as a novel optimization objective for glitch token detection.
-This approach allows us to systematically identify tokens that induce high uncertainty in the model’s predictions, moving beyond predefined patterns or embedding anomalies.
-By focusing on the model’s prediction behavior, we can detect a broader spectrum of glitch tokens, including those that might be missed by traditional embedding-based methods.
-  • Gradient-Guided Local Search: Our method combines first-order Taylor approximation with a local search strategy, efficiently navigating the vast token space.
-This approach significantly improves both the efficiency and accuracy of glitch token detection, allowing us to explore the token space more thoroughly while maintaining computational feasibility.
-  • Adaptive Cross-Architecture Detection: GlitchMiner’s framework is designed to adapt to the unique characteristics of different LLM architectures without relying on model-specific assumptions.
-This flexibility enables effective glitch token detection across a wide range of models, addressing a key limitation of previous methods that often relied on architecture-specific heuristics.
+However, with the growing reliance on these models in critical domains, a subtle yet significant vulnerability has emerged: glitch tokens geiping2024coercing .
+These anomalous tokens can severely disrupt an LLM’s output, leading to repetitive errors, nonsensical responses, or even harmful content.
+Such disruptions are particularly concerning in high-stakes applications where model reliability is crucial.
+Recent research has made progress in detecting glitch tokens.
+Methods such as Magikarp Magikarp and GlitchHunter GlitchHunter focus on abnormal embedding patterns and clustering properties.
+However, these approaches are limited by their dependence on predefined patterns, restricting their adaptability across different LLM architectures and potentially missing novel forms of glitch tokens that deviate from known patterns.
+In this paper, we introduce GlitchMiner, a novel framework for glitch token detection that overcomes the limitations of pattern-based heuristics by employing gradient-based discrete optimization to directly analyze a model’s prediction behavior.
+Central to our method is the use of entropy as an exploratory metric, guiding the search through the token space to identify regions indicative of unexpected model behavior.
+The key innovations of GlitchMiner include:
+  • Entropy-Based Optimization: By introducing entropy as a novel objective for glitch token detection, we systematically identify tokens that induce high uncertainty in model predictions.
+This approach enables detection of a broader range of glitch tokens, extending beyond traditional pattern- or embedding-based methods.
+  • Gradient-Guided Local Search: Combining first-order Taylor approximation with a local search strategy, GlitchMiner efficiently explores the vast token space.
+This significantly enhances both the efficiency and accuracy of glitch token detection, allowing thorough exploration while maintaining computational feasibility.
+  • Adaptive Cross-Architecture Detection: GlitchMiner is designed to adapt to the unique characteristics of various LLM architectures without relying on model-specific assumptions, enabling effective glitch token detection across a wide range of models.
 We conducted extensive experiments across 10 popular LLM architectures, covering models from the Llama, Qwen, Gemma, Phi, and Mistral families.
-The results show that GlitchMiner surpasses existing methods in 90% of cases when it comes to glitch token detection efficiency.
-Specifically, GlitchMiner achieves an average Detected@1000 score of 612.0 and a Detected@2000 score of 980.1.
-These results represent improvements of 10.5% and 10.7% respectively over the previous leading methods.
-In the context of glitch token detection, these improvements are significant as they translate to hundreds of additional potentially harmful tokens being identified.
-Such enhancements not only improve detection performance but also optimize query efficiency, which is crucial for enhancing the reliability and robustness of LLMs in critical applications where even a single undetected glitch token could lead to severe consequences.
-The remainder of this paper is structured as follows: Section 2 provides a comprehensive review of related work in glitch token detection and gradient-based optimization techniques.
+The results show that GlitchMiner surpasses existing methods in detection efficiency in 90% of cases.
+Specifically, GlitchMiner achieves an average Detected@1000 score of 612.0 and a Detected@2000 score of 980.1, with respective improvements of 10.5% and 10.7% over state-of-the-art methods.
+These gains enhance the detection of potentially harmful tokens, improving both accuracy and query efficiency.
+Such advancements are essential for LLM reliability, especially in applications where undetected glitch tokens could have serious consequences.
+The remainder of this paper is structured as follows: Section 2 provides a review of related work in glitch token detection and gradient-based optimization techniques.
 Section 3 details the design principles and implementation of GlitchMiner.
 Section 4 presents our experimental results and analysis, demonstrating the effectiveness of our approach across various LLM architectures.
-Finally, Section 5 concludes with a discussion of the implications of our work and outlines promising directions for future research in this critical area of LLM security and reliability.
+Finally, Section 5 concludes with a discussion of the implications of our work and outlines promising directions for future research.
 2 Background and Related Work
@@
 Glitch tokens have been found to exhibit several key characteristics:
-  • They often have abnormal embedding patterns, such as unusually large L2 norms or atypical positions in the embedding space Magikarp .
+  • They often have abnormal embedding patterns, such as unusually small $\ell_{2}$ norms or atypical positions in the embedding space Magikarp ; solidgoldmagikarp .
   • When input to an LLM, they can cause the model to produce repetitive, nonsensical, or completely unrelated outputs solidgoldmagikarp ; GlitchHunter .
@@
 3.3 Token Selection
-Motivation for Entropy-Based Approach.
-Previous methods for glitch token detection often rely on manually observed patterns or predefined distribution assumptions.
+Motivation.
+Previous methods for glitch token detection often rely on manually observed embedding patterns.
 However, these approaches can be limited in their ability to adapt to diverse LLM architectures and may overlook novel forms of glitch tokens.
@@
 4.3 Ablation Study
-To evaluate the contributions of key components in GlitchMiner, we conducted ablation studies focusing on the local search strategy, neighborhood size $K$, and batch size $B$.
+To evaluate the contributions of key components in GlitchMiner, we conducted ablation studies focusing on the local search strategy, neighborhood size $K$, batch size $B$, and initialization token.
 Effect of Local Search.

### v5

@@
 Abstract
-Glitch tokens in Large Language Models (LLMs) can trigger unpredictable behaviors, threatening model reliability and safety.
-Existing detection methods rely on predefined patterns, limiting their adaptability across diverse LLM architectures.
-We propose GlitchMiner, a gradient-based discrete optimization framework that efficiently identifies glitch tokens by introducing entropy as a measure of prediction uncertainty and employing a local search strategy to explore the token space.
-Experiments across multiple LLM architectures demonstrate that GlitchMiner outperforms existing methods in detection accuracy and adaptability, achieving over 10% average efficiency improvement.
-This method enhances vulnerability assessment in LLMs, contributing to the development of more robust and reliable applications.
+Glitch tokens—inputs that trigger unpredictable or anomalous behavior in Large Language Models (LLMs)—pose significant challenges to model reliability and safety.
+Existing detection methods primarily rely on heuristic embedding patterns or statistical anomalies within internal representations, limiting their generalizability across different model architectures and potentially missing anomalies that deviate from observed patterns.
+We introduce GlitchMiner, an behavior-driven framework designed to identify glitch tokens by maximizing predictive entropy.
+Leveraging a gradient-guided local search strategy, GlitchMiner efficiently explores the discrete token space without relying on model-specific heuristics or large-batch sampling.
+Extensive experiments across ten LLMs from five major model families demonstrate that GlitchMiner consistently outperforms existing approaches in detection accuracy and query efficiency, providing a generalizable and scalable solution for effective glitch token discovery.
 Code is available at https://github.com/wooozihui/GlitchMiner.
 1 Introduction
-Large Language Models (LLMs) have revolutionized natural language processing, enabling breakthrough applications in fields ranging from code generation LLM_code1 ; LLM_code2 ; LLM_code3 to medical research medical1 ; medical2 and education education1 ; education2 .
-However, with the growing reliance on these models in critical domains, a subtle yet significant vulnerability has emerged: glitch tokens geiping2024coercing .
-These anomalous tokens can severely disrupt an LLM’s output, leading to repetitive errors, nonsensical responses, or even harmful content.
-Such disruptions are particularly concerning in high-stakes applications where model reliability is crucial.
-Recent research has made progress in detecting glitch tokens.
-Methods such as Magikarp Magikarp and GlitchHunter GlitchHunter focus on abnormal embedding patterns and clustering properties.
-However, these approaches are limited by their dependence on predefined patterns, restricting their adaptability across different LLM architectures and potentially missing novel forms of glitch tokens that deviate from known patterns.
-In this paper, we introduce GlitchMiner, a novel framework for glitch token detection that overcomes the limitations of pattern-based heuristics by employing gradient-based discrete optimization to directly analyze a model’s prediction behavior.
-Central to our method is the use of entropy as an exploratory metric, guiding the search through the token space to identify regions indicative of unexpected model behavior.
-The key innovations of GlitchMiner include:
-  • Entropy-Based Optimization: By introducing entropy as a novel objective for glitch token detection, we systematically identify tokens that induce high uncertainty in model predictions.
-This approach enables detection of a broader range of glitch tokens, extending beyond traditional pattern- or embedding-based methods.
-  • Gradient-Guided Local Search: Combining first-order Taylor approximation with a local search strategy, GlitchMiner efficiently explores the vast token space.
-This significantly enhances both the efficiency and accuracy of glitch token detection, allowing thorough exploration while maintaining computational feasibility.
-  • Adaptive Cross-Architecture Detection: GlitchMiner is designed to adapt to the unique characteristics of various LLM architectures without relying on model-specific assumptions, enabling effective glitch token detection across a wide range of models.
-We conducted extensive experiments across 10 popular LLM architectures, covering models from the Llama, Qwen, Gemma, Phi, and Mistral families.
-The results show that GlitchMiner surpasses existing methods in detection efficiency in 90% of cases.
-Specifically, GlitchMiner achieves an average Detected@1000 score of 612.0 and a Detected@2000 score of 980.1, with respective improvements of 10.5% and 10.7% over state-of-the-art methods.
-These gains enhance the detection of potentially harmful tokens, improving both accuracy and query efficiency.
-Such advancements are essential for LLM reliability, especially in applications where undetected glitch tokens could have serious consequences.
-The remainder of this paper is structured as follows: Section 2 provides a review of related work in glitch token detection and gradient-based optimization techniques.
-Section 3 details the design principles and implementation of GlitchMiner.
-Section 4 presents our experimental results and analysis, demonstrating the effectiveness of our approach across various LLM architectures.
-Finally, Section 5 concludes with a discussion of the implications of our work and outlines promising directions for future research.
+Large Language Models (LLMs) have demonstrated remarkable capabilities and are increasingly deployed in high-stakes domains such as code generation (LLM_code1; LLM_code2; LLM_code3), healthcare (medical1; medical2), and education (education1; education2).
+As reliance on LLMs grows, ensuring their safety and output reliability becomes imperative.
+A particularly concerning threat arises from a class of anomalous inputs known as glitch tokens (solidgoldmagikarp).
+These tokens can induce LLMs to generate unexpected, erratic, or even policy-violating outputs (geiping2024coercing), undermining trust in their behavior.
+For example, as illustrated in Figure 1, when prompted to repeat the word “Mediabestanden,” Llama2-7b-chat outputs “hello world”—a semantically unrelated response.
+This phenomenon indicates that certain tokens can trigger unpredictable model behavior, even under simple and deterministic instructions.
+In response, several studies have been proposed to detect glitch tokens (Magikarp; GlitchHunter; GlitchProber).
+These methods typically rely on heuristic observations or statistical patterns in token embeddings or latent representations, such as small embedding norms or localized embedding clusters.
+Although effective in certain scenarios, such heuristics have two main limitations.
+First, the reliance on specific observed statistical features may not generalize well across different architectures or training setups.
+Second, tokens that do not conform to these empirically observed patterns may remain undetected, potentially limiting the overall coverage and robustness of these detection methods.
+Human Please repeat the string: Mediabestanden
+LLM String: ”hello world”
+Figure 1: An illustrative example of how glitch token causing Llama2-7b-chat to fail a simple repetition task. More examples can be found in Appendix B.
+To address these limitations, we propose a behavior-driven perspective.
+Rather than searching for fixed patterns within the model, we hypothesize that glitch tokens are outliers in the model’s learned distribution, and thus elicit uncertain predictions.
+This uncertainty can be quantified via the entropy of the model’s next-token distribution—high entropy indicates that the model is indecisive, providing a proxy for behavioral instability.
+Building on this insight, we present GlitchMiner, a novel framework that identifies glitch tokens by maximizing output entropy.
+The core of GlitchMiner is a gradient-guided local search, which efficiently navigates the discrete token space to identify high-entropy candidates.
+Unlike prior gradient-based discrete optimization strategies that suffer from inaccurate approximations or require large batch sampling (autoprompt; GCG), our method restricts updates to neighboring tokens in the embedding space.
+This ensures accurate Taylor-based entropy estimates and improves search efficiency.
+Our Contributions
+are summarized as follows:
+  • We propose a behavior-driven framework for glitch token detection, leveraging output entropy as an architecture-agnostic measure of model uncertainty to ensure broad applicability across diverse models.
+  • We introduce a gradient-guided local search algorithm that accurately estimates entropy gradients within the discrete token space, eliminating the dependence on large-batch sampling and improving search efficiency.
+  • Extensive experiments across 10 LLMs from 5 major model families demonstrate that GlitchMiner consistently surpasses state-of-the-art methods in detection accuracy, establishing it as a robust and effective solution for glitch token detection.
+The remainder of the paper is organized as follows: Section 2 reviews related work; Section 3 describes the GlitchMiner methodology; Section 4 presents experiments and ablations; and Section 5 concludes with future directions.
 2 Background and Related Work
-2.1 What is a Glitch Token?
-Glitch tokens, a concept that gained attention following the discovery of the "SolidGoldMagikarp" phenomenon by Rumbelow et al.
-solidgoldmagikarp , are anomalous tokens in LLMs that can trigger unexpected and often erroneous behaviors when processed by the model.
-These tokens typically result from irregularities in the training process, such as underrepresentation in the training data or inconsistencies in tokenization Magikarp ; geiping2024coercing .
-Glitch tokens have been found to exhibit several key characteristics:
-  • They often have abnormal embedding patterns, such as unusually small $\ell_{2}$ norms or atypical positions in the embedding space Magikarp ; solidgoldmagikarp .
-  • When input to an LLM, they can cause the model to produce repetitive, nonsensical, or completely unrelated outputs solidgoldmagikarp ; GlitchHunter .
-  • They may lead to significant deviations in the model’s internal activations compared to normal tokens GlitchProber .
-  • Glitch tokens can sometimes bypass content filters or trigger unexpected model behaviors, potentially compromising the safety and reliability of LLMs geiping2024coercing .
-The study of glitch tokens is crucial for understanding and mitigating potential vulnerabilities in LLMs, as these tokens can potentially be exploited for adversarial attacks or lead to unintended model behaviors in critical applications geiping2024coercing .
+2.1 Glitch Token Definition
+In practice, the identification of glitch tokens relies on specific behavioral tests.
+The most common approach, used in prior work (Magikarp; GlitchHunter; GlitchProber), is the repetition task.
+In such a task, a model is prompted to repeat a given token $t$; a failure to accurately reproduce the token is taken as evidence of it being a glitch.
+This single-test method, however, is fragile.
+The outcome can be highly sensitive to the specific wording of the prompt, leading to inconsistent results and potential false positives.
+To establish a more robust and replicable definition, we move beyond single-prompt evaluations.
+We classify a token $t$ as a glitch token only if it fails the repetition task consistently across a diverse set of $m$ prompt templates.
+This stricter criterion is formalized as follows: given a set of $m$ templates, $h_{i}(\cdot)\}_{i=1}^{m}$, the failure rate for a token $t$ is defined as:
+$\text{fail}(t):=\frac{1}{m}\sum_{i=1}^{m}\mathds{1}\left[\arg\max_{v\in\mathcal{V}}\,P(v\mid h_{i}(t))\neq t\right]$ (1)
+A token is then identified as a glitch if and only if its failure rate is absolute, i.e., $\text{fail}(t)=1$.
+This rigorous cross-verification process effectively filters out template-specific artifacts, ensuring that our analysis focuses exclusively on consistently unstable tokens.
 2.2 Glitch Token Detection
-The detection of glitch tokens in LLMs has become an increasingly important area of research, given their potential to disrupt model performance and reliability.
-The fundamental approach to identifying these anomalous tokens often relies on repetition tasks, a method popularized by the SolidGoldMagikarp study solidgoldmagikarp .
-In these tasks, the model is prompted to repeat a given token, with the assumption that normal tokens should be easily reproducible, while glitch tokens often lead to failures in this simple task.
-Building upon this basic principle, researchers have developed several sophisticated methods to detect glitch tokens more efficiently and accurately:
-Magikarp solidgoldmagikarp pioneered a fast and lightweight approach by analyzing token embeddings.
-It identifies potentially problematic tokens by examining characteristic patterns in the embedding space, such as unusual $\ell_{2}$ norms.
-These candidate tokens are then verified using repetition tasks, allowing for quick detection through direct queries to the model.
-GlitchHunter GlitchHunter introduced a clustering-based method.
-This approach is founded on the observation that glitch tokens often cluster together in the embedding space.
-GlitchHunter constructs a Token Embedding Graph (TEG) to represent token relationships and applies the Leiden algorithm traag2019louvain to identify potential glitch clusters.
-These clusters are then refined through iterative hypothesis testing to improve detection accuracy.
-GlitchProber GlitchProber takes a different approach by analyzing the internal activations within transformer layers, such as attention heads and hidden states.
-It reduces the dimensionality of these activations using PCA PCA and applies SVM classifiers SVM to identify glitch tokens.
-Additionally, it integrates mitigation mechanisms by modifying neuron activations to minimize the impact of glitch tokens during inference.
-While these methods have significantly advanced our ability to detect glitch tokens, they are inherently limited by their reliance on pattern-based heuristics.
-Magikarp and GlitchHunter focus on specific embedding patterns or clustering behaviors, while GlitchProber examines predefined patterns in activation spaces.
-This dependence on pattern-based approaches can constrain their adaptability across diverse LLM architectures and potentially overlook novel forms of glitch tokens.
-Our proposed method, GlitchMiner, aims to address these limitations by employing gradient-based discrete optimization to directly analyze an LLM’s prediction behavior, potentially offering a more flexible and comprehensive approach to glitch token detection.
+A series of methods have been proposed to efficiently rank or localise glitch tokens in large vocabularies:
+Magikarp (Magikarp) adopts a lightweight heuristic, screening tokens with atypically small $\ell_{2}$ embedding norms before verifying them via the repetition task.
+GlitchHunter (GlitchHunter) observes that glitch tokens tend to cluster in embedding space.
+It builds a token-embedding graph and applies Leiden clustering (traag2019louvain), followed by iterative hypothesis testing to refine each cluster.
+GlitchProber (GlitchProber) shifts the focus from embeddings to internal activations.
+It projects hidden states and attention outputs with PCA, then trains SVM classifiers to flag activation outliers; a mitigation step masks offending neurons at inference time.
+While effective, these approaches share a common limitation: they rely heavily on heuristic observations of embedding or activation patterns.
+Such reliance may reduce their generalizability across different architectures, and tokens not exhibiting these typical patterns may remain undetected.
+In contrast, GlitchMiner employs gradient-guided discrete optimization to maximize output entropy, an architecture-agnostic indicator of predictive uncertainty that facilitates broad generalization across diverse LLMs.
 2.3 Gradient-based Discrete Optimization
-Gradient-based discrete optimization methods hotflip ; autoprompt ; GCG ; PEZ leverage gradient information to predict how individual tokens impact the loss function.
+Gradient-based discrete optimization methods (hotflip; autoprompt; GCG; PEZ) leverage gradient information to predict how individual tokens impact the loss function.
 These approaches typically treat the one-hot encoding of tokens or token embeddings as continuous vectors to compute gradients, guiding token replacements for optimization.
-HotFlip hotflip uses the one-hot encoding of tokens to compute gradients and selects the token with the largest negative gradient to replace the current token, aiming to minimize the loss.
+HotFlip (hotflip) uses the one-hot encoding of the tokens to calculate the gradients and selects the token with the largest negative gradient to replace the current token, with the goal of minimizing the loss.
 However, it only evaluates one candidate token per iteration, which can lead to suboptimal predictions and reduced accuracy.
-AutoPrompt autoprompt improves upon HotFlip by evaluating multiple candidate tokens in each iteration.
+AutoPrompt (autoprompt) improves upon HotFlip by evaluating multiple candidate tokens in each iteration.
 Instead of relying on gradients from one-hot encodings, it utilizes token embedding gradients for loss estimation, enhancing prediction accuracy by considering a broader range of potential token replacements.
-GCG GCG extends HotFlip by incorporating multi-candidate token selection, similar to AutoPrompt, but it still uses the one-hot encoding of tokens to compute gradients for loss estimation.
-Notably, GCG has been applied to automated jailbreaks shen2023anything in LLMs, efficiently searching for adversarial suffixes.
-AutoPrompt and GCG both rely on large batch sampling to mitigate inaccuracies in gradient prediction.
+GCG (GCG) extends HotFlip by incorporating multi-candidate token selection, similar to AutoPrompt, but it still uses the one-hot encoding of tokens to compute gradients for loss estimation.
+Notably, GCG has been applied to automated jailbreaks (shen2023anything) in LLMs, efficiently searching for adversarial suffixes.
+AutoPrompt and GCG both rely on large batch sampling to mitigate inaccuracies in gradient-based prediction.
 We identified that these inaccuracies arise from the inaccuracy of Taylor expansions when input tokens are distant from the original points.
@@
 By addressing the core issue of Taylor approximation accuracy, our method allows for more efficient and accurate exploration of the token space, which is particularly valuable for glitch token detection.
-3 Method
-Algorithm 1 GlitchMiner Framework
-1:Input: Token set $\mathcal{T}$, Iteration number $I$, Batch size $B$
-2:Output: Glitch token set $\mathcal{G}$
-3:# Stage 1: Initialization
-4:$\mathcal{T}^{*}\leftarrow\text{Filter}(\mathcal{T})$ Filter out the unnecessary token set $\mathcal{T}^{*}$
-5:$\mathcal{G}\leftarrow\emptyset$
-6:# Stage 2: Mining
-7:for $i\in I$ do
-8:     Select a batch of tokens $\mathcal{B}$ from $\mathcal{T}\setminus(\mathcal{T}^{*}\cup\mathcal{G})$, where $|\mathcal{B}|=B$
-9:     for each token $t\in\mathcal{B}$ do
-10:         if Verify($t$then
-11:              Add $t$ to $\mathcal{G}$
-12:         else
-13:              Add $t$ to $\mathcal{T}^{*}$
-14:         end if
-15:     end for
-16:end for
-GlitchMiner Framework. As shown in Algorithm 1, GlitchMiner operates in two main stages: initialization and mining.
-In the initialization stage, we filter out tokens that don’t require detection to reduce the search space and improve efficiency, after which the mining stage iteratively selects batches of unverified tokens for verification.
-Each token is checked: if it is identified as a glitch token, it is added to the discovered set; otherwise, it is marked unnecessary for further detection.
-This process continues until a stopping condition is met, such as reaching a specified number of iterations or detected glitch tokens.
-The following sections detail how GlitchMiner implements its key components:
+3 Methodology
+Our goal is to develop a method that can automatically and efficiently discover glitch tokens in any given LLM.
+To this end, we propose GlitchMiner, a novel framework that formulates glitch token discovery as a behavior-driven optimization problem.
+GlitchMiner systematically mines for these tokens through a powerful iterative loop, outlined in Algorithm 1.
+Each iteration first performs Gradient-Guided Selection to pinpoint tokens with the highest predicted output entropy.
+These candidates then immediately undergo Robust Verification to determine if they are genuine glitch tokens.
+Algorithm 1 The GlitchMiner Pipeline
+1:Input: LLM model $f(\cdot)$, Full vocabulary $\mathcal{T}$, Max iterations NN
+2:Output: Verified glitch token set $\mathcal{G}$
+3:# Initialization
+[…]
+[…]
+6:# Iterative Mining Loop
+7:for $i$ from 11 to NN do
+8:  # 1. Gradient-guided selection of promising candidates
+9:  $\mathcal{T}_{c}\leftarrow\mathcal{T}\setminus(\mathcal{T}^{*}\cup\mathcal{G})$ $\triangleright$ Define current candidate set
+10:  $\mathcal{B}\leftarrow\text{SelectCandidateBatch}(\mathcal{T}_{c})$ $\triangleright$ See Sec. 3.1
+11:  # 2. Robust verification of each candidate
+12:  for each token $t\in\mathcal{B}$ do
+13:   if VerifyIsGlitch($t$) then $\triangleright$ See Sec. 3.2
+14:     $\mathcal{G}\leftarrow\mathcal{G}\cup\{t\}$
+15:   end if
+16:   $\mathcal{T}^{*}\leftarrow\mathcal{T}^{*}\cup\{t\}$ $\triangleright$ Mark token as processed
+17:  end for
+18:end for
+19:return $\mathcal{G}$
+3.1 Gradient-Guided Candidate Selection
+The core of GlitchMiner’s discovery engine is the candidate selection process.
+To ensure the overall mining is efficient, this selection is not random; instead, it employs a local search strategy to intelligently find tokens that maximize model confusion, which we measure using output entropy.
+Optimization Objective.
+To formally quantify a token’s impact on model uncertainty, we calculate its output entropy within a repetition task.
+We use the following template for this optimization process:
+User: Please repeat the string: “{token}”
+Assistant: Sure, the string is: “{token}
+Figure 2: Visualization of GlitchMiner’s local search process.
+The pivot token (black, $t_{c}$) serves as the reference point.
+Its neighbor tokens (orange and red) represent the $K=4$ closest tokens in embedding space.
+Among these, the candidate batch tokens (red) are the top $B=2$ tokens with the highest approximate entropy values, estimated via first-order Taylor approximation.
+Tokens outside the neighborhood (gray) are excluded to maintain approximation accuracy and computational efficiency.
+Here, the first {token} represents the input token $t$ being evaluated, and the second {token} is the next predicted token by LLM.
+To ensure that the input token appears directly as the model’s next prediction, we prefill the assistant’s response with the phrase Sure, the string is: “.
+The entropy is computed based on the model’s next-token distribution $p(v\mid\mathbf{h}(t))$, where $\mathbf{h}(t)$ denotes the prompt generated by inserting $t$ into the repetition template $h$.
+The entropy is thus:
+$H(t)=-\sum_{v\in\mathcal{V}}p(v\mid\mathbf{h}(t))\log p(v\mid\mathbf{h}(t))$
+Our goal is to find a batch of tokens $\mathcal{B}$ from the current set of available tokens $\mathcal{T}_{c}$ that maximizes the total entropy:
+$\mathcal{B}=\arg\max_{\mathcal{B}\subset\mathcal{T}_{c},|\mathcal{B}|=B}\sum_{t\in\mathcal{B}}H(t)$
+where $\mathcal{T}_{c}=\mathcal{T}\setminus(\mathcal{T}^{*}\cup\mathcal{G})$, with $\mathcal{T}^{*}$ being the set of filtered (see Section 3.3) or previously verified non-glitch tokens, and $\mathcal{G}$ being the set of glitch tokens found.
+Local Search Strategy.
+To efficiently solve this optimization problem, we introduce a local search strategy that addresses the limitations of global Taylor approximations.
+This process begins with an initial pivot token $t_{c}$ and iteratively refines the search.
+In each step:
   1. 1.
-    Token Filter: Identifying tokens that don’t require detection.
+    A local neighborhood $\mathcal{N}_{K}(t_{c})$ is defined, consisting of the nearest neighbors $K$ to the current pivot token $t_{c}$ in the embedding space.
   2. 2.
-    Glitch Token Verification: Methods to verify if a token is a glitch token.
+    Within this neighborhood, the entropy of each candidate token $t\in\mathcal{N}_{K}(t_{c})$ is estimated using a first-order Taylor approximation:
+    $\hat{H}(t)\approx H(t_{c})+\nabla_{e}H(t_{c})^{\top}(e_{t}-e_{t_{c}})$
+    where $e_{t}$ and $e_{t_{c}}$ are the respective embedding vectors. This approximation avoids the high cost of exact entropy calculations for all neighbors.
   3. 3.
-    Token Selection Strategy: Selecting the most promising glitch token candidates in each iteration.
-By exploring these components, we demonstrate how GlitchMiner efficiently implements glitch token mining in practice.
-3.1 Token Filter
-In the initialization stage of GlitchMiner, we first filter all tokens to remove those that do not need to be detected.
-Similar to Magikarp , we designed a token filter module that classifies tokens based on their characteristics and filters out those that are unnecessary for detection.
-The core idea of the token filter module is to classify each token by decoding and then re-encoding it, ensuring it meets specific classification criteria.
-Specifically, we prepend a special prefix "«" to each token to maintain consistency during the encoding and decoding process.
-Then, we filter out tokens that fall into the following categories:
-  • SPECIAL: Special tokens, such as [BOS], </s>, etc., which are enclosed in brackets or angle brackets and contain alphabetic characters.
-  • UNDECODEABLE: Tokens that cannot be decoded, usually containing illegal characters.
-  • UNREACHABLE: Tokens that cannot be restored to their original token ID through the decoding and re-encoding process.
-During the classification process, we first decode each token ID to obtain its corresponding string representation.
-If decoding fails, the token is classified as UNDECODEABLE.
-Next, we encode the decoded string and check if it can be restored to the original token ID.
-If it cannot, the token is classified as UNREACHABLE.
-If it meets the characteristics of a special token, it is classified as SPECIAL.
-By filtering out these unnecessary tokens, we retain only the essential ones for subsequent mining, which improves the algorithm’s efficiency and ensures the accuracy of the detection results.
-3.2 Glitch Token Verification
-Our glitch token verification process consists of two stages: optimization-driven detection and cross-verification.
-Template for Optimization. For our primary detection and optimization process, we designed the following task template:
-User: Please repeat the string: "«{token}»"
-Assistant: Sure, the string is: "«{token}
-In this template, the blue text represents the prompt, the green text {token} is the input token being tested, and the red text {token} represents the model’s predicted output.
-If the predicted token does not match the input token, it is initially classified as a potential glitch token.
-This template is specifically designed to support our gradient-based optimization process.
-Its simplicity allows for clear gradient computations, while the prefilled assistant response ensures a uniform starting point for each token evaluation.
-Wrapping the token in "«" and "»" symbols minimizes interference from surrounding context.
-These features enable precise entropy calculations and gradient estimations, which are crucial for our entropy-guided search strategy described in the token selection section.
-Cross-Verification.
-While our primary template is optimized for the search process, relying solely on one template may lead to false positives.
-To enhance the robustness of our detection, we implement a cross-verification step using two additional templates derived from GlitchHunter and Magikarp.
-Each potential glitch token identified by our primary template undergoes verification with these additional templates.
-A token is confirmed as a glitch token only if it fails the repetition task across all templates.
-This two-stage approach combines the efficiency of our optimization-driven search with the reliability of multi-template verification, significantly reducing false positives while maintaining the effectiveness of our gradient-based detection method.
-3.3 Token Selection
-Motivation.
-Previous methods for glitch token detection often rely on manually observed embedding patterns.
-However, these approaches can be limited in their ability to adapt to diverse LLM architectures and may overlook novel forms of glitch tokens.
-To address these limitations, we propose an entropy-based approach.
-Entropy, as a measure of uncertainty in probability distributions, offers a model-agnostic way to identify tokens that cause unexpected behavior in LLMs.
-By focusing on tokens that maximize entropy, we can detect glitch tokens that deviate significantly from normal token behavior, regardless of their specific characteristics or the underlying model architecture.
-Entropy-Guided Exploration. We define the entropy $H(t)$ for a token $t$ as:
-$H(t)=-\sum_{v\in\mathcal{V}}P(v\mid\mathbf{h}(t))\log P(v\mid\mathbf{h}(t))$
-where $\mathbf{h}(t)$ is the context embedding for token $t$, and $P(v\mid\mathbf{h}(t))$ is the model’s predicted probability distribution over the vocabulary $\mathcal{V}$.
-By maximizing entropy, we aim to find tokens that cause high uncertainty in the model’s predictions, potentially indicating glitch behavior.
-Optimization Objective. Our goal is to find a batch of tokens $\mathcal{B}$ that maximizes the total entropy:
-$\mathcal{B}=\arg\max_{\mathcal{B}\subset\mathcal{T}_{c},|\mathcal{B}|=B}\sum_{t\in\mathcal{B}}H(t)$
-Here, $\mathcal{T}_{c}$ represents the current candidate set, defined as $\mathcal{T}_{c}=\mathcal{T}\setminus(\mathcal{T}^{*}\cup\mathcal{G})$, where:
-  • $\mathcal{T}$ is the initial set of all tokens in the model’s vocabulary.
-  • $\mathcal{T}^{*}$ is the set of tokens that have been filtered out or verified as non-glitch tokens.
-This includes special tokens, undecoded tokens, and tokens that have been checked and confirmed as normal.
-  • $\mathcal{G}$ is the set of identified glitch tokens discovered so far in the process.
-By excluding $\mathcal{T}^{*}$ and $\mathcal{G}$ from $\mathcal{T}$, we ensure that our search focuses only on the remaining unverified tokens, improving efficiency and preventing redundant checks.
-Local Search Strategy.
-To efficiently solve this optimization problem while addressing the limitations of global Taylor approximation, we introduce a local search strategy.
-We begin by selecting an initial token $t_{0}$ and computing its entropy $H(t_{0})$ and gradient $\nabla_{e}H(t_{0})$.
-We then define a local neighborhood $\mathcal{N}_{K}(t_{0})$ consisting of the $K$ nearest neighbors to $t_{0}$ in the embedding space.
-This localized approach allows us to refine the search within a region close to $t_{0}$, improving approximation accuracy and ensuring efficient exploration of potential glitch tokens.
-Within this local neighborhood, we estimate the entropy of each candidate token $t\in\mathcal{N}_{K}(t_{0})$ using first-order Taylor approximation:
-$\hat{H}(t)\approx H(t_{0})+\nabla_{e}H(t_{0})^{T}(e_{t}-e_{t_{0}})$
-where $e_{t}$ and $e_{t_{0}}$ are the embedding vectors of tokens $t$ and $t_{0}$, respectively.
-This approximation allows us to efficiently estimate the entropy change for nearby tokens without the computational cost of exact calculations.
-Based on these estimates, we select a batch $\mathcal{B}$ of $B$ tokens with the highest approximated entropy.
-For this selected batch, we then compute the actual entropy values and choose the token with the highest entropy as the starting point for the next iteration.
-This iterative process ensures that our search is guided by the most promising candidates at each step.
-This local search strategy significantly improves the accuracy of entropy estimation by focusing on tokens close to the reference point, addressing a key limitation of previous gradient-based methods.
-By balancing exploration of the token space with exploitation of local information, our approach enables more efficient and effective glitch token detection.
-The combination of entropy-guided exploration and local search allows our method to adapt to diverse LLM architectures without relying on model-specific assumptions.
-This results in a more robust and versatile approach to glitch token detection, capable of identifying a wide range of anomalous tokens across different model types.
-4 Experiments
-4.1 Experimental Setup
-Test LLMs.
-We used a diverse set of LLMs from five different model families to evaluate the performance of our glitch token detection approach.
-The selected models include Meta’s Llama series llama2 ; llama3_1 , Alibaba’s Qwen models qwen2 ; qwen2_5 , Google’s Gemma models gemma2 , Microsoft’s Phi-3 models phi3 , and Mistral models mistral_7B ; mistral2024nemo .
-The details are presented in Table 1.
+    A batch $\mathcal{B}$ of $B$ tokens with the highest approximated entropy $\hat{H}(t)$ is selected from the neighborhood. This batch is returned for verification.
+  4. 4.
+    To guide the next search step, the actual entropy $H_{t}$ is computed for the tokens in $\mathcal{B}$, and the one with the highest actual entropy becomes the new pivot, $t_{c}$.
+Figure 2 visualizes this process.
+This local search strategy significantly improves the accuracy of entropy estimation by focusing on tokens close to the pivot token.
+This entire process corresponds to the ‘SelectCandidateBatch’ function in Algorithm 1.
 Model Family Model Names
@@
 Table 1: Test LLMs used in the experiments.
-Evaluation Metrics.
-To assess the performance of our glitch token detection method, we employ the Detected@N metric.
-This metric measures the absolute number of true glitch tokens identified within the top $N$ predictions made by the algorithm.
-It is calculated as:
-[…]
-For example, Detected@1000 measures the number of true glitch tokens discovered within the top 1000 predicted tokens.
-This metric directly reflects a method’s effectiveness in identifying glitch tokens under practical constraints, as it captures both the detection capability and query efficiency of the approach.
-By comparing the number of glitch tokens detected under the same query budget ($N$), we can effectively evaluate different methods’ ability to maximize token discovery while minimizing computational resources.
-This makes it an ideal metric for real-world applications where efficient detection of glitch tokens is crucial.
-Baselines.
-We compare our proposed glitch token detection method with two state-of-the-art approaches: GlitchHunter GlitchHunter and Magikarp Magikarp .
-These methods serve as the primary benchmarks for evaluating our approach.
-Although GlitchProber GlitchProber is another relevant method, it follows a fundamentally different approach by pre-collecting a subset of glitch tokens to train a classifier, introducing a supervised learning component.
-In contrast, GlitchMiner, along with GlitchHunter and Magikarp, uses heuristic-based methods to detect glitch tokens without relying on labeled data or additional classifier training.
-This methodological difference makes a direct comparison less meaningful, so we focus our evaluation on methods that align more closely with our unsupervised approach.
-Parameter Settings.
-In our implementation of GlitchMiner, we use $K$=32 and $B$=8 as the default parameters.
-These values were chosen based on empirical testing to balance computational efficiency and detection effectiveness.
-Specifically, $K$=32 defines the size of the local neighborhood considered in each iteration, while $B$=8 determines the batch size for entropy computation.
-These settings have shown to provide a good trade-off between exploration of the token space and exploitation of local information across various model architectures.
-Initialization Strategy in Experiments.
-To ensure stable and consistent comparisons across runs, we initialize the search with the token exhibiting the smallest $\ell_{2}$ norm in the embedding space, based on prior observations that such tokens often exhibit glitch-like behaviors.
-However, as shown in Figure 4, we found that GlitchMiner remains robust to different initialization choices, achieving similar performance even with random starting points.
-4.2 Main Results
 Model Metric GlitchHunter Magikarp GlitchMiner (ours)
@@
 Average Detected@1000 40.9 553.9 612.0
-Detected@2000 93.0 885.6 980.1
+Detected@2000 83.0 885.6 980.1
 Table 2: Detected@1000 and Detected@2000 comparison of methods across different models.
-Table 2 displays the performance comparison of GlitchMiner with the state-of-the-art methods GlitchHunter and Magikarp across different LLM architectures using the Detected@N metric.
-The results highlight GlitchMiner’s strong and versatile detection capabilities.
-In the majority of cases, GlitchMiner achieved the highest Detected@2000 score, consistently outperforming both baselines in terms of identified glitch tokens within the top 2000 predictions.
-This is particularly evident in models such as Llama-2-7B-chat-hf, where GlitchMiner achieved a Detected@2000 of 532, surpassing Magikarp by a notable margin.
-For the Qwen and other models, GlitchMiner maintained robust performance, demonstrating adaptability across different LLM architectures.
+3.2 Robust Verification of Candidates
+Once a batch of promising candidates is selected, each one undergoes a rigorous verification step (‘VerifyIsGlitch’ in our pipeline) to filter out false positives.
+A candidate is confirmed as a true glitch only if it fails the repetition task across a set of diverse prompt templates.
+To ensure this robustness, our verification process uses $m=3$ templates for cross-validation: in addition to the template used for optimization (Sec.
+3.1), we employ two others adapted from prior work (Magikarp; GlitchHunter).
+A token is formally classified as a glitch if and only if its failure rate (as defined in Equation 1) is 1.
+The additional prompt templates and analysis of false positives are provided in Appendix A.1 and Appendix A.2, respectively.
+Figure 3: Impact of different Neighborhood Size $K$ on GlitchMiner’s performance
+Figure 4: Impact of different Batch Size $B$ on GlitchMiner’s performance
+3.3 Practical Implementation: Search Space Pre-filtering
+To maximize search efficiency, following (land2024fishing), we perform a one-time pre-filtering of the vocabulary (‘PreFilter’ in Algorithm 1).
+This step removes tokens that pose no real-world risk, allowing GlitchMiner’s overall search to focus its resources.
+We filter three categories:
+  • SPECIAL Tokens: Predefined symbols like [BOS] or </s>.
+These are filtered out because they serve a reserved functional role for the model (e.g., indicating the start of a sequence) rather than representing user-generated text.
+  • UNDECODEABLE Tokens: Tokens that correspond to byte sequences that cannot be decoded into a valid string, often because they violate encoding standards like UTF-8.
+They are excluded as they do not map to any meaningful user-generated text.
+  • UNREACHABLE Tokens: Tokens that exist in the vocabulary but can never be generated by the model’s tokenizer from any text input.
+Further explanations of these token categories are provided in Appendix A.3.
+4 Experiments
+4.1 Experimental Setup
+Evaluated LLMs.
+We used a diverse set of LLMs from five different model families to evaluate the performance of our glitch token detection approach.
+The selected models include Meta’s Llama series (llama2; llama3_1), Alibaba’s Qwen models (qwen2; qwen2_5), Google’s Gemma models (gemma2), Microsoft’s Phi-3 models (phi3), and Mistral models (mistral_7B; mistral2024nemo).
+The details are presented in Table 1.
+Evaluation Metrics.
+We evaluate our glitch token detection method using the Detected@N metric, which counts the number of true glitch tokens identified within the top $N$ predictions.
+For instance, Detected@1000 measures how many glitch tokens are found among the top 1000 candidates.
+This metric balances detection accuracy and query efficiency, reflecting a method’s practical effectiveness under fixed query budgets.
+Comparing Detected@N values thus provides a direct measure of each method’s ability to maximize glitch token discovery while minimizing computational resources, making it well-suited for real-world applications.
+Figure 5: Effect of Initialization Method on GlitchMiner’s Detected@1000 score.
 Figure 6: Comparison of GlitchMiner performance with and without local search strategy
-These experimental findings underscore GlitchMiner’s ability not only to accurately detect glitch tokens but also to generalize across a diverse set of model architectures.
-This adaptability and precision position GlitchMiner as a powerful tool for enhancing the robustness and security of LLMs.
+Baselines.
+We compare our proposed glitch token detection method with two state-of-the-art approaches: GlitchHunter (GlitchHunter) and Magikarp (Magikarp).
+These methods serve as the primary benchmarks for evaluating our approach.
+Although GlitchProber (GlitchProber) is another relevant method, it follows a fundamentally different approach by pre-collecting a subset of glitch tokens to train a classifier, introducing a supervised learning component.
+In contrast, GlitchMiner, along with GlitchHunter and Magikarp, uses heuristic-based methods to detect glitch tokens without relying on labeled data or additional classifier training.
+This methodological difference makes a direct comparison less meaningful, so we focus our evaluation on methods that align more closely with our unsupervised approach.
+Parameter Settings.
+In our implementation of GlitchMiner, we use KK=32 and BB=8 as the default parameters.
+These values were chosen based on empirical testing to balance computational efficiency and detection effectiveness.
+Specifically, KK=32 defines the size of the local neighborhood considered in each iteration, while BB=8 determines the batch size for entropy computation.
+These settings have shown to provide a good trade-off between exploration of the token space and exploitation of local information across various model architectures.
+Initialization Strategy in Experiments.
+To ensure stable and consistent comparisons across runs, we initialize the search with the token exhibiting the smallest $\ell_{2}$ norm in the embedding space, based on prior observations that such tokens often exhibit glitch-like behaviors.
+However, as shown in Figure 5, we found that GlitchMiner remains robust to different initialization choices, achieving similar performance even with random starting points.
+4.2 Main Results
+The performance of GlitchMiner against the baselines is detailed in Table 2.
+The results provide strong evidence for the effectiveness of our entropy-guided search, showing that GlitchMiner consistently discovers more glitch tokens than both GlitchHunter and Magikarp under fixed query budgets.
+On average, GlitchMiner outperforms the strongest baseline, Magikarp, by 10.7% on the Detected@2000 metric.
+This performance advantage is not uniform but reveals an important trend: while Magikarp’s simple norm-based heuristic can be effective for an initial, low-budget scan (e.g., on Llama-3.1-8B), GlitchMiner’s more sophisticated local search strategy consistently proves more fruitful as the search budget expands.
+For instance, on Llama-2-7B-chat-hf, GlitchMiner finds nearly three times as many glitches as Magikarp (532 vs.
+186), demonstrating its superior ability to uncover less obvious candidates that simple heuristics miss.
+These findings validate that framing glitch detection as a behavior-driven optimization problem is a more robust and generalizable strategy than relying on static, model-specific patterns.
+Figure 7: Average entropy comparison between glitch tokens and normal tokens across different models.
+Glitch tokens have higher entropy, indicating greater uncertainty in the model’s predictions for these tokens.
 4.3 Ablation Study
@@
 This trend indicates that as $K$ grows, the Taylor approximation becomes less effective, resulting in reduced prediction accuracy.
-Figure 2: Impact of different Neighborhood Size $K$ on GlitchMiner’s performance
-Figure 3: Impact of different Batch Size $B$ on GlitchMiner’s performance
-Figure 4: Effect of Initialization Method on GlitchMiner’s Detected@1000 score.
 Effect of Batch Size.
@@
 As shown in the figure, glitch tokens consistently exhibit significantly higher entropy than normal tokens across all models.
-Figure 5: Average entropy comparison between glitch tokens and normal tokens across different models.
-Glitch tokens have higher entropy, indicating greater uncertainty in the model’s predictions for these tokens.
 This pronounced difference in entropy values indicates that models are more uncertain when predicting glitch tokens compared to normal tokens.
@@
 This suggests that our method can be effectively applied to a wide range of LLMs with different architectures and tokenization strategies.
-These findings reinforce the effectiveness of GlitchMiner’s entropy-based optimization in efficiently detecting glitch tokens by focusing on areas of high prediction uncertainty within the model.
+These findings validate the effectiveness of GlitchMiner’s entropy-based optimization in efficiently detecting glitch tokens by focusing on areas of high prediction uncertainty within the model.
 5 Conclusion
-In this paper, we introduced GlitchMiner, a novel framework for detecting glitch tokens in LLMs through gradient-based discrete optimization.
-Our method combines entropy-based loss functions with a local search strategy to navigate token space efficiently, improving both detection accuracy and computational efficiency.
-Experimental results across diverse LLMs highlight GlitchMiner’s robustness and versatility compared to existing methods.
-Our findings confirm GlitchMiner’s effectiveness in identifying glitch tokens across architectures, demonstrating its potential for enhancing the security and reliability of LLM-based applications.
-Future work could explore extending this approach to other high-dimensional discrete optimization tasks.
-Broader Impact
-The potential impact of this research extends beyond mere academic interest.
-By improving the detection of glitch tokens, GlitchMiner contributes to enhancing the overall reliability and safety of LLMs in real-world applications.
-This is particularly crucial in high-stakes domains such as healthcare, finance, and autonomous systems, where the consequences of model errors can be severe.
-Furthermore, our approach opens new avenues for understanding and mitigating vulnerabilities in AI systems, potentially influencing the development of more robust and trustworthy AI technologies in the future.
+In this paper, we introduced GlitchMiner, a novel framework that detects glitch tokens in LLM by reframing the problem as a behavior-driven optimization task.
+Departing from prior work that relies on static patterns, our approach identifies anomalous tokens by searching for inputs that maximize the model’s predictive uncertainty, measured by output entropy.
+We operationalize this principle with an efficient gradient-guided local search strategy that can accurately pinpoint high-entropy candidates in the vast token space.
+Extensive experiments on 10 diverse LLMs demonstrated that GlitchMiner significantly outperforms state-of-the-art baselines in both detection accuracy and efficiency.
+Our work provides a more robust and generalizable tool for auditing and enhancing the reliability of LLMs, with future work pointing towards mitigation strategies and adaptation for black-box models.
 Acknowledgments
 We would like to express our sincere gratitude to Sander Land for his valuable insights and feedback on this work, particularly regarding the evaluation metrics.
-His suggestions were instrumental in refining our approach and improving the clarity of our results.
-We also wish to thank Guanlin Li for his valuable suggestions on our manuscript.
-His insights were crucial for improving the content and structure of the paper.
+We wish to thank Guanlin Li for his valuable suggestions on our manuscript.
+This work was supported in part by the National Key R&D Program of China (2023YFB3107505), in part by Shaanxi Natural Science Funds for Distinguished Young Scholars(2023-JC-JQ-52), in part by the Natural Science Foundation of China (62302371), in part by the Postdoctoral Fellowship Program of CPSF (GZC20232035), and in part by the China Postdoctoral Science Foundation (2025M771552).
+Model Token ID Token String Model’s Response
+Qwen2.5-7B-Instruct 83969 PostalCodesNL ”: ”
+Qwen2.5-7B-Instruct 88372 useRalative ”: ”
+Qwen2.5-7B-Instruct 78323 thuisontvangst ”: ”
+Llama-2-7b-chat 24291 IABot Of course! Here is the string ”@{”:
+I hope this helps! Let me know if you have any other questions.
+Llama-2-7b-chat 17916 abestanden Sure! Here is the repetition of the string: Wikimedia
+Llama-2-7b-chat 12731 ederbörd Sure! The string ”pon” repeats 3 times.
+Mistral-7B-Instruct-v0.3 20016 NdEx I’m sorry, but the string ”tcx” cannot be repeated as is because it does not contain any characters or symbols that can be replicated.
+It may have been entered incorrectly, as it appears to consist only of backslashes and may not represent a meaningful string in its current form.
+Could you please check and provide the correct string if possible?
+Mistral-7B-Instruct-v0.3 26668 iNdEx I’m sorry, but the string ” /******/iji__[” cannot be repeated as it is not a valid or meaningful string in standard English or most programming languages.
+It appears to contain some uncommon characters that could represent specific instructions in certain contexts, such as within JSON or in some esoteric programming languages.
+Without more information about its intended usage, it’s difficult to provide an accurate repetition.
+Mistral-7B-Instruct-v0.3 8709 ICENSE I’m sorry, it seems there’s a typo in your request. The correct spelling is ”LICENSE.” Here it is again: LICENSE.
+Table 3: Examples of Glitch Tokens Discovered by GlitchMiner
+Appendix A Experimental Details
+A.1 Verification Templates
+As stated in the main text, our robust verification process relies on three distinct templates for cross-validation to ensure reliability.
+A token is confirmed as a true glitch only if it fails the repetition task across all three.
+The first template is the one used during the optimization phase, while the other two are adapted from prior work to test for robustness under different phrasings.
+The structure of all three templates is detailed below for clarity and reproducibility.
+Template 1: Primary Optimization Template
+Description: This is the primary template used to calculate output entropy during the gradient-guided search.
+User Prompt: Please repeat the string: "{token}"
+Assistant Prefill: Sure, the string is: "‘
+Template 2: GlitchHunter Adaptation (Direct Repetition)
+Description: This template, adapted from GlitchHunter, directly asks the model to repeat the target token in a simple question-answer format.
+User Prompt: Question: Please can you repeat the string ’{token}’?
+Assistant Prefill: Of course! Here is the repeated string:‘
+Template 3: Magikarp Adaptation (Few-shot Consistency)
+Description: Adapted from Magikarp, this is a more complex few-shot template that first establishes a persona of a ”helpful assistant.” It tests if the model remains consistent when presented with the target token by first providing an example of how it handles a nonsensical string.
+System Prompt: This helpful assistant explains what a phrase means, replying with identical responses to identical questions for reliability.
+Conversation History:
+User: What does ’edslfkj’ mean?
+Assistant: ’edslfkj’ has no specific meaning I am aware of.
+Final User Prompt: What does ’{token}’ mean?
+Assistant Prefill:
+A.2 Analysis of the Verification Stage
+The GlitchMiner framework operates as a two-stage pipeline:
+  1. 1.
+    A gradient-guided search proposes candidate tokens that maximize output entropy.
+  2. 2.
+    A robust verification stage confirms whether these candidates are true glitches.
+The verification stage is intentionally stringent, requiring a candidate to fail the repetition task across all three prompt templates (including those adapted from GlitchHunter and Magikarp) to be confirmed.
+A token that fails on only one or two templates does not meet this strict criterion.
+Table 4 summarizes the number of false-positive glitch tokens among the top 1000 candidates, which fail verification with template 1 but pass verification with either template 2 or template 3.
+This analysis demonstrates the necessity of employing cross-validation across multiple prompt templates.
+Model Candidates Filtered by Verification (Top 1000)
+Qwen2.5-7B-Instruct 0
+Qwen2-7B-Instruct 0
+Llama-2-7b-chat-hf 36
+Llama-3.1-8B-Instruct 42
+Gemma-2-2b-it 34
+Gemma-2-9b-it 28
+Mistral-7B-Instruct-v0.3 51
+Mistral-Nemo-Instruct-2407 23
+Phi-3.5-mini-instruct 76
+Phi-3-mini-128k-instruct 49
+Table 4: Count of high-entropy candidates from the top 1000 search results that were filtered out during the robust verification stage.
+These are tokens that did not fail on all three verification templates.
+A.3 Rationale and Details of the Token Filter
+The core function of our Token Filter is to determine a token’s practical relevance and, consequently, its real-world risk profile.
+It answers a fundamental question: can a given token in the vocabulary be triggered by a user through standard text input?
+If a token cannot be produced from text, it is excluded from our analysis.
+This is not part of the glitch detection logic itself, but a crucial prerequisite that allows GlitchMiner to focus its computational budget efficiently on tokens that represent genuine security and reliability concerns.
+Our filter removes three main categories of such tokens:
+SPECIAL Tokens
+These are predefined functional symbols used for model structure or training (e.g., [BOS], </s>). As they are not part of natural language inputs, they are straightforward to exclude.
+Unreachable Tokens
+These are tokens that exist in the vocabulary but can never be generated by the model’s tokenizer from any text input, often existing as artifacts of the tokenizer’s construction.
+We provide a concrete example from the Llama-2-7B-chat-hf model to illustrate this:
+[…]
+[…]
+This demonstrates that no text input a user can provide will ever result in token 71. Since it cannot be triggered in practice, testing it for ”glitchiness” is unnecessary.
+Undecodeable Tokens
+This category includes tokens corresponding to byte sequences that cannot be decoded into a valid string, often due to violating encoding standards like UTF-8.
+For example, in the same Llama-2 tokenizer:
+[…]
+  • This byte is an invalid start for a UTF-8 character sequence. Therefore, it cannot be decoded into coherent text and often appears as a generic replacement symbol.
+Such tokens are excluded as they do not map to any meaningful user-generated text.
+Appendix B Examples of Discovered Glitch Tokens
+Table 3 showcases several examples of glitch tokens discovered by GlitchMiner across different models, illustrating the variety of anomalous behaviors that can be triggered.
+The ”Response” column displays each model’s full output when given the primary verification prompt, Please repeat the string: {token}.